### v1

# Generalized dessin partition function, cut-and-join type representation, and polynomiality

Abstract.
In this paper we study the enumeration problem of a special kind of branched coverings of the Riemann sphere, whose partition function can be regarded as a generalization of the dessin partition function. First we derive a cut-and-join type representation for the generalized dessin partition function using the actions of $W$-type operators on Schur functions. Then we study the polynomial structures of the generalized dessin partition function. We prove a polynomial structure for the disconnected correlators which generalizes a result of Stanley, and then propose a conjectural polynomial structure for the connected correlators. We prove the $1$-point case of this conjecture using the KP integrability of this partition function.
1. Introduction
The enumeration of branched coverings of Riemann surfaces is one of the most important subjects in the study of modern mathematical physics. Fix the number of branch points and the ramification type over each branch point, then the weighted counting of isomorphism classes of such branched coverings are the so-called Hurwitz numbers [21]. Hurwitz numbers are known to be closely related to many other mathematical theories, such as representation theory, integrable hierarchies, combinatorics, and the geometry of moduli spaces of algebraic curves and Gromov-Witten theory, see e.g. [8, 13, 16, 17, 19, 26, 23, 30].
Among various types of branched coverings of Riemann surfaces, the branched coverings of ${\mathbb{C}}{\mathbb{P}}^{1}$ over three points $0,1,\infty\in{\mathbb{C}}{\mathbb{P}}^{1}$ are of particular interest since it leads to the notion of Grothendieck’s dessin d’enfants [19, 5], which are bi-colored ribbon graphs on surfaces. For an introduction to Grothendieck’s dessin d’enfants, ribbon graphs, and some related topics, see the book [26]. The enumeration of dessin d’enfants has been well-studied by mathematical physicists, see e.g. [2, 3, 11, 24, 25, 34, 36, 37, 38]. Let the monodromy type over $\infty\in{\mathbb{C}}{\mathbb{P}}^{1}$ and the numbers of the preimages of 00 and $1$ be fixed, then the generating series of the weighted counting of such branched coverings over ${\mathbb{C}}{\mathbb{P}}^{1}$ is called the dessin partition function by Kazarian and Zograf in [24]. In their paper, they have derived the Virasoro constraints and Eynard-Orantin topological recursion for the dessin partition function using the combinatorial properties of ribbon graphs. As a consequence of the Virasoro constraints, they obtain a cut-and-join representation for the dessin partition function which naturally leads to the KP integrability.
In this work, we study a generalization of the dessin partition function. Fix a positive integer $r$, and consider the branched coverings over ${\mathbb{C}}{\mathbb{P}}^{1}$ with $r+1$ branch points $q_{1},\cdots,q_{r},q_{r+1}\in{\mathbb{C}}{\mathbb{P}}^{1}$, such that the preimage $f^{-1}(q_{i})$ consists of $k_{i}$ points for $1\leq i\leq r$, and the ramification type over $q_{r+1}$ is described by a partition of integer $\mu=(\mu_{1},\cdots,\mu_{l})$. Denote by
$N_{k_{1},\cdots,k_{r}}^{\bullet}(\mu)=\sum_{f}\frac{1}{|\operatorname{Aut}(f)|}$
the weighted counting of equivalence classes of such branched coverings from a (not necessarily connected) Riemann surface to ${\mathbb{C}}{\mathbb{P}}^{1}$, and denote by $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$ the generating series:
(1) $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=\sum_{k_{1},\cdots,k_{r}=1}^{\infty}\sum_{\mu\in{\mathcal{P}}}\Big{(}\prod_{i=1}^{r}v_{i}^{k_{i}}\Big{)}s^{|\mu|}N_{k_{1},\cdots,k_{r}}^{\bullet}(\mu)\cdot p_{\mu},$
where $s$, $\bm{v}=(v_{1},\cdots,v_{r})$ and $\bm{p}=(p_{1},p_{2},\cdots)$ are some formal variables, and $p_{\mu}=p_{\mu_{1}}p_{\mu_{2}}\cdots p_{\mu_{l}}$ for a partition $\mu=(\mu_{1},\cdots,\mu_{l})$. We will call $Z_{(r)}(s;\bm{v};\bm{p})$ the generalized dessin partition function. When $r=2$, it becomes the dessin partition function studied in [3, 24, 38, 36, 37, 34]. We are mainly concerned with the following three aspects of this partition function:
  • Cut-and-join type representation;
  • KP integrability;
  • Polynomiality of the correlators.
In particular, we will derive a cut-and-join type representation for $Z_{(r)}$ in this work. The KP integrability for $Z_{(r)}$ follows immediately from the cut-and-join type representation; and we make an attempt to study the polynomiality of the connected correlators (i.e., the coefficients of $\log Z_{(r)}(s;\bm{v};\bm{p})$) using the KP integrability. There are some other strategies to prove the polynomiality of various types of Hurwitz numbers in literatures, for examples, by relating this problem to the intersection theory over moduli spaces, by analyzing the cut-and-join equation, and by using certain fermionic operators and combinatorics. See e.g. [13, 12, 17, 18, 25] and references therein for exhaustive study.
Now we briefly state the main results of this work. Let $P_{-1}^{(k)}$ be the cut-and-join type operators (see §3.1 for a review of the notations):
(2) $P_{-1}^{(k)}=\operatorname{Res}_{z=0}\Big{(}:z^{k-2}\big{(}\partial_{z}+J(z)\big{)}^{k-1}J(z):\Big{)}dz,$
where $J(z)=\sum_{n\in{\mathbb{Z}}_{+}}nt_{n}\cdot z^{n-1}+\sum_{n\in{\mathbb{Z}}_{+}}\frac{\partial}{\partial t_{n}}z^{-n-1}$ is the generating function of Heisenberg operators. Our first main theorem is the following cut-and-join type representation of the generalized dessin partition function (see Theorem 3.2 and Proposition 3.2):
Theorem 1.1.
Let $\bm{v}=(v_{1},v_{2},\cdots,v_{r})$, and let $a_{1}(\bm{v}),\cdots,a_{r+1}(\bm{v})$ be the polynomials in $\bm{v}$ given by:
(3) $a_{k}(\bm{v})=\frac{(-1)^{k-1}}{k!}e_{r}(\bm{v})+\frac{1}{k}\sum_{j=0}^{r}\Big{(}\sum_{i=1}^{k-1}\frac{(-1)^{i+k-1}\cdot i^{r-j}}{i!\cdot(k-1-i)!}\Big{)}e_{j}(\bm{v}),$
where $e_{j}(\bm{v})$ is the elementary symmetric polynomial of degree $j$ in $v_{1},v_{2},\cdots,v_{r}$ for $1\leq j\leq r$, and $e_{0}(\bm{v})=1$. Then there is a cut-and-join type representation for the generalized dessin partition function $Z(s;v_{1},\cdots,v_{r};\bm{p})$:
(4) $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=\exp\Big{(}s\cdot\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}\Big{)}(1).$
When $r=2$, this theorem recovers the cut-and-join representation for the dessin partition function derived by Kazarian-Zograf in [24]. It is worth noting that Kazarian and Zograf proved the cut-and-join representation using the combinatorics of ribbon graphs. Here we prove Theorem 1.1 by a different combinatorial method, namely, the actions of $W$-type operators on Schur functions [27], which enables us to treat $Z_{(r)}$ for general $r\geq 1$ in a unified fashion. The general $r$ case was also studied in [2, 3] using the matrix model. Alexandrov-Mironov-Morozov-Natanzon [2] derived another cut-and-join type representation for this partition function using infinite Casimir operators. By contrast, we use only finite many $W-$type operators for each partition function.
In literatures it is known that the above cut-and-join type operators $P_{-1}^{(k)}$ belong to the algebra $W_{1+\infty}$ which gives the symmetry of the KP hierarchy [14], and thus the partition function $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$ is a tau-functions of the KP hierarchy for every $r$. Using the Schur function expansion of $Z_{(r)}$, we are able to write down the affine coordinates of this tau-function on the Sato-Grassmannian explicitly:
(5) $\begin{split}a_{n,m}^{(r)}=(-1)^{n}\cdot\frac{s^{m+n+1}}{(m+n+1)\cdot m!\cdot n!}\cdot\prod_{i=1}^{r}\prod_{j=-n}^{m}(v_{i}+j).\end{split}$
When $r=2$, this recovers the affine coordinates for the dessin partition function computed by Zhou in [36]. Now one can use a formula of Zhou [35, §5] to compute the connected correlators of the tau-function $Z_{(r)}$ using the affine coordinates (5). For details and concrete examples, see §3.5.
Then we study the polynomiality of the correlators of the tau-function $Z_{(r)}$. The following polynomial structure for the disconnected correlators is a straightforward generalization of a result of Stanley [33], see §4 for details:
Theorem 1.2.
Fix two partitions of integers $\lambda=(\lambda_{1},\cdots,\lambda_{k})$ (where $k\leq r$) and $\mu=(\mu_{1},\cdots,\mu_{l})$. Then there is a polynomial $A_{(r)}^{\lambda,\mu}(x)\in{\mathbb{Q}}[x]$, such that the disconnected correlators of $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$ satisfies:
(6) $n!\cdot N_{n-\lambda_{1},\cdots,n-\lambda_{k},n,\cdots,n}^{\bullet}\big{(}(\mu_{1},\cdots,\mu_{l},1^{n-|\mu|})\big{)}=A_{(r)}^{\lambda,\mu}(n)$
for every positive integer $n$ with $n\geq|\mu|$.
About the connected correlators $N_{k_{1},\cdots,k_{r}}^{\circ}(\mu)$, we propose the following conjecture about their polynomial structure (see §5 for details):
Conjecture 1.1.
Fix some positive integers $l$ and $k_{1},k_{2},\cdots,k_{r}$. There exists a polynomial $A_{k_{1},\cdots,k_{r}}(x_{1},\cdots x_{l})\in{\mathbb{Q}}[x_{1},\cdots,x_{l}]$ such that
(7) $A_{k_{1},\cdots,k_{r}}(\mu_{1},\cdots\mu_{l})=z_{\mu}\cdot N_{|\mu|-k_{1},|\mu|-k_{2},\cdots,|\mu|-k_{r-1},k_{r}}^{\circ}(\mu)$
for every partition $\mu=(\mu_{1},\cdots,\mu_{l})$ satisfying $|\mu|>\max\{k_{1},\cdots,k_{r}\}$.
This kind of polynomiality, for various kinds of Hurwitz numbers, was widely studied in literatures (see e.g. [13, 9, 12, 16, 17, 18, 25] and references therein) and related to many important problems in mathematical physics. Based on numerical data, we can also propose a stronger version of this conjecture, which enables certain coefficients to be non-negative integers. See Remark 5.1 for more details.
Since one of our goals is to explore the relation between KP integrability and the polynomiality, we present a proof of the $1$-point case (i.e., the case $l=1$) using the affine coordinates (5) and Zhou’s formula [35]. The computations for general $l>1$ become rather complicated.
The rest of this paper is arranged as follows. In §2 we give a brief review of the preliminaries about Hurwitz numbers. In §3 we derive the Schur function expansion, cut-and-join type representation, and KP integrability of the partition function $Z_{(r)}$. In §4 we use Stanley’s method to prove Theorem 1.2. Finally in §5, we present some examples of Conjecture 1.1 and prove the $1$-point case using the KP integrability of $Z_{(r)}$.
2. Hurwitz Numbers, Burnside Formula, and Dessins d’Enfants
First we recall some preliminaries of Schur functions, Hurwitz numbers, constellations, and Grothendieck’s dessins d’enfants. The materials in this section can be found in the books of Macdonald [29], Cavalieri-Miles [6], and Lando-Zvonkine [26].
2.1. Partitions of integers and Schur functions
First we recall the notions of partitions of integers and Schur functions. For reference, see [29, §I].
A partition of a nonnegative integer $n$ is a sequence of nonnegative integers $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{l})$ satisfying $\lambda_{1}\geq\cdots\geq\lambda_{l}>0$ and $|\lambda|=\lambda_{1}+\cdots+\lambda_{l}=n$. In particular, the empty partition $\lambda=(\emptyset)$ is a partition of 00. The number $l$ is called the length of this partition and will be denoted by $l(\lambda)$. Let $\lambda=(\lambda_{1},\cdots\lambda_{l})$ be a partition, and denote by $m_{i}(\lambda)$ the number of $i$ appearing in the sequence $(\lambda_{1},\cdots\lambda_{l})$. We will use the following notation:
(8) $z_{\lambda}=\prod_{i\geq 1}i^{m_{i}(\lambda)}\cdot m_{i}(\lambda)!.$
There is a one-to-one correspondence between the set of partitions of integers and the set of Young diagrams, The Young diagram corresponding to $\lambda=(\lambda_{1},\cdots,\lambda_{l(\lambda)})$ consists of $|\lambda|$ boxes, such that there are exactly $\lambda_{i}$ boxes in the $i$-th row. Let $\lambda=(\lambda_{1},\cdots,\lambda_{l(\lambda)})$ be a partition, then the transpose $\lambda^{t}=(\lambda_{1}^{t},\cdots,\lambda_{m}^{t})$ of $\lambda$ is the partition given by $m=\lambda_{1}$ and $\lambda_{j}^{t}=|\{i|\lambda_{i}\geq j\}|=\sum_{k\geq j}m_{k}(\lambda)$. The Young diagram associated with $\lambda^{t}$ can be obtained from the Young diagram associated with $\lambda$ by flipping along the diagonal. It is clear that $(\lambda^{t})^{t}=\lambda$ for every partition $\lambda$. The Frobenius notation of a partition $\lambda$ is defined to be:
$\lambda=(m_{1},m_{2},\cdots,m_{k}|n_{1},n_{2},\cdots,n_{k}),$
where $k$ is the number of boxes on the diagonal of the Young diagram, and
$m_{i}=\lambda_{i}-i,\qquad n_{i}=\lambda_{i}^{t}-i,\qquad i=1,2,\cdots,k.$
Let $\lambda=(\lambda_{1},\cdots,\lambda_{l(\lambda)})$ be a partition of integer. Let $\square$ be a box in the Young diagram associated with $\lambda$, and in such a case we will use the notation $\square\in\lambda$. If $\square$ is at the $i$-th row and $j$-th column of the Young diagram, then the hook length of the box $\square$ is defined to be:
(9) $h(\square)=(\lambda_{i}-i)+(\lambda_{j}^{t}-j)+1,$
and the content $c(\square)$ is defined to be:
(10) $c(\square)=j-i.$
Now we recall the definition of the Schur function $s_{\lambda}$ indexed by a partition $\lambda$. First consider the special case $\lambda=(m|n)=(m+1,1^{n})$. The Young diagram associated with such a partition is a hook, and in this case $s_{(m|n)}$ is defined by:
$s_{(m|n)}=h_{m+1}e_{n}-h_{m+2}e_{n-1}+\cdots+(-1)^{n}h_{m+n+1},$
where $h_{n}$ and $e_{n}$ are the complete symmetric function and elementary symmetric function of degree $n$ respectively. For a general $\lambda=(m_{1},\cdots,m_{k}|n_{1},\cdots,n_{k})$, the Schur function $s_{\lambda}$ is defined by:
$s_{\lambda}=\det(s_{(m_{i}|n_{j})})_{1\leq i,j\leq k}.$
There are also equivalent definitions for $s_{\lambda}$:
$s_{\lambda}=\det(h_{\lambda_{i}-i+j})_{1\leq i,j\leq n}=\det(e_{\lambda_{i}^{t}-i+j})_{1\leq i,j\leq m},$
for $n\geq l(\lambda)$ and $m\geq l(\lambda^{t})$. The Schur function indexed by the empty partition is defined to be $s_{(\emptyset)}=1$. Denote by $\Lambda$ the space of all symmetric functions, then the set of all Schur functions $\{s_{\lambda}\}_{\lambda}$ is a basis of $\Lambda$.
It is well-known that $\{p_{\lambda}\}$ gives another basis for $\Lambda$, where $p_{\lambda}=p_{\lambda_{1}}\cdots p_{\lambda_{l}}$ for a partition $\lambda=(\lambda_{1},\cdots,\lambda_{l})$ and $p_{k}$ is the Newton symmetric function of degree $k$. The above two bases for $\Lambda$ are related by (see e.g. [29]):
(11) $s_{\mu}=\sum_{\lambda}\frac{\chi^{\mu}(C_{\lambda})}{z_{\lambda}}p_{\lambda},$
where $\chi^{\mu}$ is the irreducible character of the symmetric group $S_{|\mu|}$ indexed by the partition $\mu$, and $C_{\lambda}\subset S_{|\mu|}$ is the conjugacy class whose cycle type is $\lambda$. The summations on the right-hand sides of (11) are over all partitions $\lambda$ with $|\lambda|=|\mu|$. For an introduction of the representation theory of $S_{n}$, see e.g. [15, §4].
In some literatures (see e.g. [27]), the authors regard $s_{\mu}$ as a function in the variables $\bm{t}=(t_{1},t_{2},\cdots)$ via (11) where $t_{n}=\frac{p_{n}}{n}$, and call it the Schur polynomial indexed by $\mu$. In fact, if one defines a degree by setting $\deg(t_{n})=n$, then it is easy to see that $s_{\mu}(\bm{t})$ is a homogeneous polynomial of degree $|\mu|$ in $t_{1},t_{2},\cdots,t_{|\mu|}$. In this paper we will use such a notation.
2.2. Hurwitz numbers labeled by $m$ partitions
In this subsection we recall the definition of Hurwitz numbers. Here we only consider branched covers of the Riemann sphere ${\mathbb{C}}{\mathbb{P}}^{1}$.
Let $d$ be a positive integer, and let $\mu^{(1)},\mu^{(2)}\cdots,\mu^{(m)}$ be $m$ partitions of $d$. Let $f:\Sigma_{g}\to{\mathbb{C}}{\mathbb{P}}^{1}$ be a branched covering of the Riemann sphere by a (not necessarily connected) Riemann surface $\Sigma_{g}$ of genus $g$, such that there are $m$ branch points $q_{1},\cdots,q_{m}\in{\mathbb{C}}{\mathbb{P}}^{1}$, and the ramification type of $f$ over $q_{i}$ is $\mu^{(i)}$. More precisely, the pre-image of the point $q_{i}$ consists of $l(\mu^{(i)})$ distinct points $p_{1},\cdots,p_{l(\mu^{(i)})}$ on $\Sigma_{g}$, such that $f$ locally looks like $z\mapsto z^{\mu^{(i)}_{j}}$ in some local coordinate $z$ near $p_{j}$. The Riemann-Hurwitz formula tells that:
(12) $2g-2=d(m-2)-\sum_{i=1}^{m}l(\mu^{(i)}).$
The (disconnected) Hurwitz number $H_{g}^{\bullet}(\mu^{(1)},\cdots,\mu^{(m)})$ is defined to be the weighted counting of all such branched coverings $f$:
$H_{g}^{\bullet}(\mu^{(1)},\cdots,\mu^{(m)})=\sum_{f}\frac{1}{|\operatorname{Aut}(f)|},$
where $\operatorname{Aut}(f)$ is the group of automorphisms of the branched covering $f$. Similarly, one can define the connected Hurwitz number $H_{g}^{\bullet}(\mu^{(1)},\cdots,\mu^{(m)})$ by
$H_{g}^{\circ}(\mu^{(1)},\cdots,\mu^{(m)})=\sum_{f:\text{ connected}}\frac{1}{|\operatorname{Aut}(f)|},$
where the summation is over branched coverings $f$ of ramification type $\mu^{(1)},\cdots,\mu^{(m)}$ from a connected Riemann surface $\Sigma_{g}$.
2.3. Group-theoretic description of Hurwitz numbers
There is a group-theoretic description of Hurwitz numbers, which leads naturally to the concept of constellations.
Let $f:\Sigma_{g}\to{\mathbb{C}}{\mathbb{P}}^{1}$ be a branched covering of degree $d$ with $m$ branch points $q_{1},q_{2},\cdots,q_{m}\in{\mathbb{C}}{\mathbb{P}}^{1}$, such that the ramification type over $q_{i}$ is described by a partition $\mu^{(i)}$. Choose a point $x_{0}\in{\mathbb{C}}{\mathbb{P}}^{1}\backslash\{q_{1},\cdots,q_{m}\}$, then the map $f$ induces a group homomorphism (see e.g. [6, §7] for details)
$\Phi:\pi_{1}({\mathbb{C}}{\mathbb{P}}^{1}\backslash\{q_{1},\cdots,q_{m}\})\to S_{d},$
called the monodromy representation, where $S_{d}$ is the permutation group of the points in the pre-image $f^{-1}(x_{0})$ of $x$, and
$\pi_{1}({\mathbb{C}}{\mathbb{P}}^{1}\backslash\{q_{1},\cdots,q_{m}\})=\langle\gamma_{1},\cdots,\gamma_{m}|\gamma_{1}\cdots\gamma_{m}=e\rangle.$
Here $\gamma_{i}$ denotes the homotopy class of a small loop around $q_{i}$, and the cycle type of the permutation $\Phi(\gamma_{i})$ is given by the partition $\mu^{(i)}$. The monodromy representation is called connected if its image $Im(\Phi)$ acts transitively on $\{1,2,\cdots,d\}$. The connectedness of a branched covering $f$ is equivalent to the connectedness of the associated monodromy representation.
Conversely, one can also construct a branched covering $f$ of ${\mathbb{C}}{\mathbb{P}}^{1}$ (up to equivalence) from a given monodromy representation. Thus one obtains a bijection
$\{[f:\Sigma_{g}\to{\mathbb{C}}{\mathbb{P}}^{1}]\}\to\operatorname{Hom}\big{(}\pi_{1}({\mathbb{C}}{\mathbb{P}}^{1}\backslash\{q_{1},\cdots,q_{m}\}),S_{d}\big{)}/\sim,$
where $[f]$ denotes the equivalence classes of $f$, and $\sim$ is the conjugation by elements in $S_{d}$. Then the Hurwitz numbers can also be defined algebraically as (see [6, 26]):
(13) $H_{g}^{\bullet}(\mu^{(1)},\cdots,\mu^{(m)})=\frac{1}{d!}\big{|}\{(\alpha_{1},\cdots,\alpha_{m})\in(S_{d})^{m}|\alpha_{1}\cdots\alpha_{m}=e,\alpha_{i}\in C_{i}\}\big{|},$
where $d=|\mu^{(i)}|$, and $C_{i}\subset S_{d}$ is the conjugacy class whose cycle type is given by $\mu^{(i)}$. By definition, the Hurwitz numbers is equal to zero unless $g$ and $\mu^{(i)}$ satisfy the Riemann-Hurwitz formula (12). And similarly,
(14) $\begin{split}H_{g}^{\circ}(\mu^{(1)},\cdots,\mu^{(m)})=\frac{1}{d!}\big{|}\{(\alpha_{1},\cdots,&\alpha_{m})\in(S_{d})^{m}|\alpha_{1}\cdots\alpha_{m}=e,\alpha_{i}\in C_{i},\\ &\text{and $\langle\alpha_{1},\cdots,\alpha_{m}\rangle$ acts transitively}\}\big{|}.\end{split}$
The above construction leads naturally to the notion of constellations, see e.g. [26, §1] for reference. A $k$-constellation is a sequence $[\sigma_{1},\cdots,\sigma_{k}]$ of permutations $\sigma_{i}\in S_{d}$, such that:
  • 1)
    $\sigma_{1}\sigma_{2}\cdots\sigma_{k}=e\in S_{d}$;
  • 2)
    the subgroup $\langle\sigma_{1},\cdots,\sigma_{k}\rangle$ acts transitively on the set $\{1,\cdots,d\}$.
Here $k$ and $d$ are called the length and degree of this constellation respectively. Then the computation of connected Hurwitz numbers of genus $g$ and ramification type $(\mu^{(1)},\cdots,\mu^{(k)})$, is actually equivalent to the enumeration of constellations $[\sigma_{1},\cdots,\sigma_{k}]$ of length $k$ and degree $d$, such that the cycle type of each $\sigma_{i}$ is exactly $\mu^{(i)}$. The genus $g$ is determined from $\mu^{(i)}$ by the Riemann-Hurwitz formula (12).
2.4. Burnside formula for Hurwitz numbers
One can compute the disconnected Hurwitz numbers algebraically using the above group-theoretic description. Let $C_{i}\subset S_{d}$ be the conjugacy class in $S_{d}$ whose cycle type is given by the partition $\mu^{(i)}$, then the following famous result is called the Burnside character formula (see e.g. [6, §9] for a proof):
(15) $H_{g}^{\bullet}(\mu^{(1)},\cdots,\mu^{(m)})=\sum_{|\eta|=d}\Big{(}\frac{\dim V^{\eta}}{d!}\Big{)}^{2}\cdot\prod_{i=1}^{m}\frac{|C_{i}|\cdot\chi^{\eta}(C_{i})}{\dim V^{\eta}},$
where $V^{\eta}$ denotes the irreducible representation of $S_{d}$ indexed by the partition $\eta$, and $\chi^{\eta}$ is the irreducible character associated with $V^{\eta}$. The summation in the right-hand side is over all partitions $\eta$ of $d$.
The numbers $|C_{i}|$, $\chi^{\eta}(C_{i})$, and $\dim V^{\eta}$ appearing in the right-hand side of the above formula can all be represented combinatorially in terms of the partitions $\eta$ and $\mu^{(i)}$. The size of the conjugacy class $C_{i}$ is:
(16) $|C_{i}|=\frac{d!}{\prod_{k\geq 1}k^{m_{k}}\cdot m_{k}(\mu^{(i)})!}=\frac{d!}{z_{\mu^{(i)}}},$
and the dimension of $V^{\eta}$ is given by the hook length formula:
(17) $\dim V^{\eta}=\frac{d!}{\prod_{\square\in\eta}h(\square)},$
where $\square\in\eta$ means $\square$ runs over boxes in the Young diagram associated with $\eta$, and $h(\square)$ is the hook length of the box $\square$. The evaluation of the character $\chi^{\eta}(C_{i})$ can be computed using the Frobenius formula, and for details we refer to [15, §4].
2.5. Enumeration of Grothendieck’s dessins d’enfants
In this subsection, we briefly review the enumeration of Grothendieck’s dessins d’enfants.
Let $X$ be a smooth algebraic curve of genus $g$ defined over ${\mathbb{C}}$. A famous result of Belyĭ [5] tells that $X$ can be defined over the field of algebraic numbers $\overline{\mathbb{Q}}$ if and only if there exists a holomorphic map $f:X\to{\mathbb{C}}{\mathbb{P}}^{1}$ such that $f$ is ramified only over three points $0,1,\infty\in{\mathbb{C}}{\mathbb{P}}^{1}$. Here $(X,f)$ is called a Belyĭ pair if such a branched covering $f$ exists. Given a Belyĭ pair $(X,f)$, consider the pre-image $f^{-1}([0,1])$ of the segment $[0,1]\subset{\mathbb{C}}{\mathbb{P}}^{1}$ in the curve $X$. Denote by $d$ the degree of the branched covering $f$, then the pre-image $f^{-1}([0,1])$ is a ribbon graph of genus $g$ with $d$ edges, whose vertices are bicolored such that adjacent vertices have different colors. Here the vertices of this ribbon graph are actually the pre-image of the two points 00 and $1$, and the vertices in $f^{-1}(0)$ and $f^{-1}(1)$ are distinguished by two colors. In [19], Grothendieck shows that isomorphism classes of Belyĭ pairs $(X,f)$ are in one-to-one correspondence with such connected bicolored ribbon graphs, and these connected bicolored ribbon graphs are called Grothendieck’s dessins d’enfants.
The enumeration of Grothendieck’s dessins d’enfants has been discussed in literatures of mathematical physics, see e.g. [2, 3, 11, 24, 25, 34, 36, 37, 38]. Let $\mu=(\mu_{1},\cdots,\mu_{m})$ be a partition of integer, and consider the weighted counting of dessins:
$N_{k,l}(\mu)=\sum_{\Gamma\in{\mathcal{D}}_{k,l;\mu}}\frac{1}{|\operatorname{Aut}(\Gamma)|},$
where $k,l$ are two positive integers, and ${\mathcal{D}}_{k,l;\mu}$ is the set of dessins whose associated Belyĭ pair $(X,f)$ is of degree $d=|\mu|$ and satisfies:
  • 1)
    $|f^{-1}(0)|=k$, and $|f^{-1}(1)|=l$;
  • 2)
    The ramification type over $\infty\in{\mathbb{C}}{\mathbb{P}}^{1}$ is described by the partition $\mu$,
and $\operatorname{Aut}(\Gamma)$ denotes the group of automorphisms of the dessin $\Gamma$ that preserves each boundary component. The above dessin counting number $N_{k,l}(\mu)$ is actually a summation of Hurwitz numbers which count the branched coverings of ${\mathbb{C}}{\mathbb{P}}^{1}$ ramified over three points, such that the preimage of the first two points consists of $k$ and $l$ elements respectively, and the ramification type over the third point is described by $\mu$. In [24], Kazarian and Zograf derived the Virasoro constraints, Eynard-Orantin topological recursion, and a cut-and-join equation for this dessin counting problem using combinatorial properties of the dessins d’enfants, and proved that the generating series of $N_{k,l}(\mu)$ is the logarithm of a tau-function of the KP hierarchy. The KP-affine coordinates and emergent geometry of this tau-function has been discussed by Zhou in [36, 37].
It is natural to generalize the above dessin counting problem to the case where the number of branch points on ${\mathbb{C}}{\mathbb{P}}^{1}$ is not necessarily to be three. Let $\mu=(\mu_{1},\cdots,\mu_{l})$ be a partition of integer, and let $1\leq k_{1},\cdots,k_{r}$ be a sequence of positive integers (where $r\geq 1$). Consider the branched coverings $f:\Sigma_{g}\to{\mathbb{C}}{\mathbb{P}}^{1}$ from a connected Riemann surface to ${\mathbb{C}}{\mathbb{P}}^{1}$ ramified over $r+1$ points $q_{1},\cdots,q_{r},q_{r+1}\in{\mathbb{C}}{\mathbb{P}}^{1}$, such that the preimage $f^{-1}(q_{i})$ consists of $k_{i}$ elements for $1\leq i\leq r$, and the ramification type over $q_{r+1}$ is described by $\mu$. We denote by
(18) $N_{k_{1},\cdots,k_{r}}^{\circ}(\mu)=\sum_{f:\text{ connected}}\frac{1}{|\operatorname{Aut}(f)|}$
the weighted counting of all such branched coverings. Similarly, we denote by
(19) $N_{k_{1},\cdots,k_{r}}^{\bullet}(\mu)=\sum_{f}\frac{1}{|\operatorname{Aut}(f)|}$
the weighted counting of branched coverings of this ramification type where we allow the Riemann surface $\Sigma_{g}$ to be disconnected. It is clear that $N_{k_{1},\cdots,k_{r}}^{\bullet}(\mu)=N_{k_{1},\cdots,k_{r}}^{\circ}(\mu)=0$ unless $k_{i}\leq|\mu|$ for every $i$.
Now denote by ${\mathcal{P}}$ the set of all partitions of integers (here we allow the empty partition $(\emptyset)\in{\mathcal{P}}$). In this paper, we study the following generalization of the dessin partition function:
(20) $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=\sum_{k_{1},\cdots,k_{r}=1}^{\infty}\sum_{\mu\in{\mathcal{P}}}\Big{(}\prod_{i=1}^{r}v_{i}^{k_{i}}\Big{)}s^{|\mu|}N_{k_{1},\cdots,k_{r}}^{\bullet}(\mu)\cdot p_{\mu},$
where $s,v_{1},\cdots,v_{r}$ and $\bm{p}=(p_{1},p_{2},\cdots)$ are formal variables, and we denote
$p_{\mu}=p_{\mu_{1}}p_{\mu_{2}}\cdots p_{\mu_{l}}$
for a partition $\mu=(\mu_{1},\mu_{2},\cdots,\mu_{l})$. In the special case $r=2$, the weighted counting $N_{k_{1},k_{2}}^{\circ}(\mu)$ recovers the dessin counting $N_{k_{1},k_{2}}(\mu)$, and the partition function $Z_{(2)}(s;v_{1},v_{2};\bm{p})$ coincide with the dessin partition function studied in [24].
3. Cut-and-Join Representation for the Generalized Dessin Partition Functions
In this section, we prove Theorem 1.1. In order to derive this cut-and-join type representation for the generalized dessin partition function $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$, we need to represent this partition function in terms of Schur functions first, and then apply the results in [27]. As a straightforward consequence of the cut-and-join type representation, one easily sees that $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$ is a tau-function of the KP hierarchy. We will also write down the affine coordinates of this tau-function, and compute the connected $n$-point functions using Zhou’s formula, which will be useful in §5 when studying the polynomiality.
3.1. Actions of $W$-type operators on Schur functions
In this subsection we recall the actions of $W$-type operators on Schur functions derived by X. Liu and the second-named author in [27], which will be our main tool for deriving the cut-and-join type representation in §3.3.
Let $k\in{\mathbb{Z}}_{+}$, and denote:
(21) $P^{(k)}(z)=:\Big{(}\partial_{z}+J(z)\Big{)}^{k-1}J(z):,$
where ::::$\quad$:: : is the normal ordering of bosonic operators (see e.g. [7, §5] for the definition of the normal ordering) and $J(z)$ is the following operator:
(22) $J(z)=\sum_{n\in{\mathbb{Z}}_{+}}nt_{n}\cdot z^{n-1}+\sum_{n\in{\mathbb{Z}}_{+}}\frac{\partial}{\partial t_{n}}z^{-n-1}.$
Define the operators $\{P_{m}^{(k)}\}_{m\in{\mathbb{Z}}}$ by
(23) $P^{(k)}(z)=\sum_{m\in{\mathbb{Z}}}P_{m}^{(k)}\cdot z^{-m-k},$
then the actions of these operators on Schur functions are given by:
Theorem 3.1 ([27]).
Let $\mu=(\mu_{1},\cdots,\mu_{l})$ be a partition of integer. Then for every $k\in{\mathbb{Z}}_{+}$ and $m\in{\mathbb{Z}}$, one has:
(24) $\begin{split}P_{m}^{(k)}s_{\mu}(\bm{t})=&\sum_{i=1}^{l}k[\mu_{i}-m-i]_{k-1}\cdot s_{\mu-m\epsilon_{i}}(\bm{t})+\delta_{m,0}[-l]_{k}\cdot s_{\mu}(\bm{t})\\ &+\delta_{m<0}\cdot\sum_{n=1}^{-m}(-1)^{m-n}k[n-l-1]_{k-1}\cdot s_{(\mu,n,1^{-m-n})}(\bm{t}),\end{split}$
where $\mu-m\epsilon_{i}=(\mu_{1},\cdots,\mu_{i-1},\mu_{i}-m,\mu_{i+1},\cdots,\mu_{l})$ (in the case where this sequence is no longer a partition, see §2 in [27] for more details), and
(25) $[a]_{k}=\begin{cases}\prod_{j=0}^{k-1}(a-j),\qquad&k\geq 1;\\ 1,&k=0,\end{cases}$
for an integer $a$. Here we use the conventions $1^{n}=(1,\cdots,1)\in{\mathbb{Z}}^{n}$ and $1^{0}=\emptyset$.
3.2. Schur function expansion for the generalized dessin partition functions
Now we use the Burnside formula and Lemma 3.1 to derive a Schur function expansion for the generalized dessin partition function $Z(s;v_{1},\cdots,v_{r};\bm{p})$.
Let $\mu$ be a partition of the integer $d$, and denote by ${\mathcal{P}}_{d}$ the set of all partitions of $d$. By the Burnside character formula (15) one has:
$\begin{split}N_{k_{1},\cdots,k_{r}}^{\bullet}(\mu)=&\sum_{\begin{subarray}{c}\mu^{(1)},\cdots,\mu^{(r)}\in{\mathcal{P}}_{d}\\ \text{s.t. }l(\mu^{(i)})=k_{i}\end{subarray}}\sum_{\eta\in{\mathcal{P}}_{d}}\Big{(}\frac{\dim V^{\eta}}{d!}\Big{)}^{2}\cdot\frac{|C|\cdot\chi^{\eta}(C)}{\dim V^{\eta}}\prod_{i=1}^{r}\frac{|C_{i}|\cdot\chi^{\eta}(C_{i})}{\dim V^{\eta}}\\ =&\sum_{\begin{subarray}{c}\mu^{(1)},\cdots,\mu^{(r)}\in{\mathcal{P}}_{d}\\ \text{s.t. }l(\mu^{(i)})=k_{i}\end{subarray}}\sum_{\eta\in{\mathcal{P}}_{d}}\Big{(}\frac{\dim V^{\eta}}{d!}\Big{)}^{2}\cdot\frac{d!\cdot\chi^{\eta}(C)}{z_{\mu}\cdot\dim V^{\eta}}\prod_{i=1}^{r}\frac{d!\cdot\chi^{\eta}(C_{i})}{z_{\mu^{(i)}}\cdot\dim V^{\eta}}\\ =&\sum_{\begin{subarray}{c}\mu^{(1)},\cdots,\mu^{(r)}\in{\mathcal{P}}_{d}\\ \text{s.t. }l(\mu^{(i)})=k_{i}\end{subarray}}\sum_{\eta\in{\mathcal{P}}_{d}}\Big{(}\frac{d!}{\dim V^{\eta}}\Big{)}^{r-1}\cdot\frac{\chi^{\eta}(C)}{z_{\mu}}\prod_{i=1}^{r}\frac{\chi^{\eta}(C_{i})}{z_{\mu^{(i)}}},\end{split}$
where $C_{i}\subset S_{d}$ (resp. $C\subset S_{d}$) is the conjugacy class whose cycle type is $\mu^{(i)}$ (resp. $\mu$). And then by (11):
$\begin{split}&Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})\\ =&\sum_{d=0}^{\infty}\sum_{\eta,\mu,\mu^{(1)},\cdots,\mu^{(r)}\in{\mathcal{P}}_{d}}\Big{(}\frac{d!}{\dim V^{\eta}}\Big{)}^{r-1}\cdot\frac{\chi^{\eta}(C)}{z_{\mu}}s^{d}p_{\mu}\prod_{i=1}^{r}\frac{\chi^{\eta}(C_{i})\cdot v_{i}^{l(\mu^{(i)})}}{z_{\mu^{(i)}}}\\ =&\sum_{d=0}^{\infty}\sum_{\eta,\mu^{(1)},\cdots,\mu^{(r)}\in{\mathcal{P}}_{d}}\Big{(}\frac{d!}{\dim V^{\eta}}\Big{)}^{r-1}\cdot s^{d}\cdot s_{\eta}(\bm{t})\prod_{i=1}^{r}\frac{\chi^{\eta}(C_{i})\cdot v_{i}^{l(\mu^{(i)})}}{z_{\mu^{(i)}}}\\ \end{split}$
where $\bm{t}=(t_{1},t_{2},t_{3},\cdots)$, and
(26) $t_{n}=\frac{p_{n}}{n},\qquad n\geq 1.$
For each fixed $i$ ($1\leq i\leq r$), taking $p_{k}=v_{i}$ for every $k\geq 1$ in (11) gives:
$\sum_{\mu^{(i)}\in{\mathcal{P}}}\frac{\chi^{\eta}(C_{i})\cdot v_{i}^{l(\mu^{(i)})}}{z_{\mu^{(i)}}}=s_{\eta}(\bm{t})\big{|}_{p_{1}=p_{2}=p_{3}=\cdots=v_{i}}=s_{\eta}(t_{k}=\frac{v_{i}}{k}),$
and thus we have:
(27) $\begin{split}Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=&\sum_{d=0}^{\infty}\sum_{\eta\in{\mathcal{P}}_{d}}\Big{(}\frac{d!}{\dim V^{\eta}}\Big{)}^{r-1}\cdot s^{d}\cdot s_{\eta}(\bm{t})\prod_{i=1}^{r}s_{\eta}(t_{k}=\frac{v_{i}}{k})\\ =&\sum_{\eta\in{\mathcal{P}}}\frac{s^{|\eta|}\cdot(|\eta|!)^{r-1}}{(\dim V^{\eta})^{r-1}}\cdot s_{\eta}(\bm{t})\prod_{i=1}^{r}s_{\eta}(t_{k}=\frac{v_{i}}{k}).\end{split}$
The following evaluation formula is well-known in literatures (see e.g. [29, §I.3. Example 4]; and see Appendix A for a new proof using Theorem 3.1):
Lemma 3.1.
Take $t_{k}=\frac{v}{k}$ for every $k\geq 1$, then one has:
(28) $s_{\mu}(t_{k}=\frac{v}{k})=\prod_{\square\in\mu}\frac{v+c(\square)}{h(\square)},$
where $c(\square)$ and $h(\square)$ are the content and hook-length respectively of the box $\square\in\mu$ in the Young diagram associated with $\mu$.
Now apply Lemma 3.1 to (27), and we obtain the following:
$Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=\sum_{\eta\in{\mathcal{P}}}\frac{s^{|\eta|}\cdot(|\eta|!)^{r-1}}{(\dim V^{\eta})^{r-1}}\cdot s_{\eta}(\bm{t})\prod_{i=1}^{r}\prod_{\square\in\eta}\frac{v_{i}+c(\square)}{h(\square)}.$
Recall that $\dim V^{\eta}$ is given by (17), and thus:
(29) $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=\sum_{\eta\in{\mathcal{P}}}s^{|\eta|}\cdot\Big{(}\prod_{\square\in\eta}\frac{\prod_{i=1}^{r}\big{(}v_{i}+c(\square)\big{)}}{h(\square)}\Big{)}\cdot s_{\eta}(\bm{t}).$
One can further simplify this formula by using the following well-known evaluation formula for Schur functions (see e.g. [29, §I.3. Example 5]):
(30) $s_{\eta}(\delta_{k,1})=\frac{1}{\prod_{\square\in\eta}h(\square)},$
where $s_{\eta}(\delta_{k,1})$ is the evaluation of the Schur polynomial $s_{\eta}(\bm{t})$ at $t_{k}=\delta_{k,1}$ for every $k\geq 1$. In this way we can conclude that:
Proposition 3.1.
Take $t_{n}=\frac{p_{n}}{n}$, then there is a Schur function expansion for the generalized dessin partition function:
(31) $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=\sum_{\eta\in{\mathcal{P}}}s^{|\eta|}\cdot\Big{(}\prod_{\square\in\eta}\prod_{i=1}^{r}\big{(}v_{i}+c(\square)\big{)}\Big{)}\cdot s_{\eta}(\delta_{k,1})s_{\eta}(\bm{t}).$
Remark 3.1.
In the special case $r=2$, the above result recovers the Schur function expansion for the dessin partition function (cf. [36, (5)]):
(32) $Z_{\text{dessin}}=\sum_{\mu\in{\mathcal{P}}}s_{\mu}\cdot\prod_{\square\in\mu}\frac{s\cdot\big{(}u+c(\square)\big{)}\big{(}v+c(\square)\big{)}}{h(\square)},$
where $u=v_{1}$ and $v=v_{2}$.
3.3. Cut-and-join type representation for the generalized dessin partition functions
Now we are ready to derive the cut-and-join type representation (or $W$-type representation) for the generalized dessin partition functions $Z(s;v_{1},\cdots,v_{r};\bm{p})$. Fix an integer $r\geq 1$, and in this subsection we prove the following:
Theorem 3.2.
Let $\bm{v}=(v_{1},v_{2},\cdots,v_{r})$, and let $a_{1}(\bm{v}),\cdots,a_{r+1}(\bm{v})$ be the polynomials in $\bm{v}$ determined by:
(33) $\sum_{k=1}^{r+1}ka_{k}(\bm{v})\cdot\prod_{j=0}^{k-2}(x-j)=\prod_{i=1}^{r}(x+v_{i}).$
Then there is a cut-and-join type representation for the generalized dessin partition function $Z(s;v_{1},\cdots,v_{r};\bm{p})$:
(34) $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=e^{s\cdot\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}}(1),$
where $P_{-1}^{(k)}$ are the W-type operators defined by (21)-(23).
Proof.
Take $m=-1$ in Theorem 3.1, and we obtain:
(35) $P_{-1}^{(k)}s_{\mu}(\bm{t})=\sum_{i=1}^{l(\mu)}k[\mu_{i}+1-i]_{k-1}\cdot s_{\mu+\epsilon_{i}}(\bm{t})+k[-l(\mu)]_{k-1}\cdot s_{(\mu,1)}(\bm{t}).$
The right-hand side is actually a summation over all possible ways of adding a box on the right or at the bottom of the Young diagram associated with $\mu$ such that the resulting graph still represents a partition. We denote by $\square$ this new box, and denote by $\mu+\square$ the resulting partition. If the new box $\square$ is in the $i$-th row where $1\leq i\leq l(\mu)$, then its content is $c(\square)=\mu_{i}+1-i$; and if $\square$ is in the $(l(\mu)+1)$-th row, its content is $c(\square)=-l(\mu)$. Thus (35) becomes:
$\begin{split}P_{-1}^{(k)}s_{\mu}(\bm{t})=&\sum_{\mu+\square\in{\mathcal{P}}}k[c(\square)]_{k-1}\cdot s_{\mu+\square}(\bm{t})\\ =&\sum_{\mu+\square\in{\mathcal{P}}}k\cdot\prod_{j=0}^{k-2}\big{(}c(\square)-j\big{)}\cdot s_{\mu+\square}(\bm{t}).\end{split}$
Then we have:
$\begin{split}\Big{(}\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}\Big{)}s_{\mu}=&\sum_{k=1}^{r+1}a_{k}(\bm{v})\sum_{\mu+\square\in{\mathcal{P}}}k\prod_{j=0}^{k-2}\big{(}c(\square)-j\big{)}s_{\mu+\square}(\bm{t})\\ =&\sum_{\mu+\square\in{\mathcal{P}}}\Big{(}\sum_{k=1}^{r+1}ka_{k}(\bm{v})\prod_{j=0}^{k-2}\big{(}c(\square)-j\big{)}\Big{)}s_{\mu+\square}(\bm{t})\\ =&\sum_{\mu+\square\in{\mathcal{P}}}\Big{(}\prod_{i=1}^{r}\big{(}c(\square)+v_{i}\big{)}\Big{)}\cdot s_{\mu+\square}(\bm{t}),\end{split}$
and thus:
$\begin{split}e^{s\cdot\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}}(1)=&e^{s\cdot\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}}\big{(}s_{(\emptyset)}(\bm{t})\big{)}\\ =&\sum_{n=0}^{\infty}\frac{s^{n}}{n!}\cdot\Big{(}\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}\Big{)}^{n}\big{(}s_{(\emptyset)}(\bm{t})\big{)}\\ =&\sum_{\mu\in{\mathcal{P}}}\frac{s^{|\mu|}}{|\mu|!}\cdot K_{\mu}\cdot\prod_{\square\in\mu}\prod_{i=1}^{r}\big{(}c(\square)+v_{i}\big{)}s_{\mu}(\bm{t}),\end{split}$
where $K_{\mu}$ is the number of all possible ways to add $n$ boxes successively to ()($\emptyset$)( ∅ ) to obtain the Young diagram of $\mu$, i.e., the number of standard Young tableaux of shape $\mu$. The number of standard Young tableaux of shape $\mu$ is known to be equal to the dimension of the irreducible representation $V^{\mu}$ (see e.g. [15, §4.3]):
(36) $K_{\mu}=\dim V^{\mu}=\frac{|\mu|!}{\prod_{\square\in\mu}h(\square)},$
And thus:
(37) $\begin{split}e^{s\cdot\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}}(1)=&\sum_{\mu\in{\mathcal{P}}}s^{|\mu|}\cdot\prod_{\square\in\mu}\frac{\prod_{i=1}^{r}\big{(}c(\square)+v_{i}\big{)}}{h(\square)}s_{\mu}(\bm{t}).\end{split}$
Then (34) follows from this equality and (29). ∎
As a straightforward consequence of the above theorem, we have the following:
Corollary 3.1.
The generalized dessin partition function $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$ satisfies the cut-and-join type equation:
(38) $\frac{\partial}{\partial s}Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=\sum_{k=1}^{r+1}a_{k}(\bm{v})\cdot P_{-1}^{(k)}Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p}).$
Example 3.1.
First consider the simplest case $r=1$. Here (33) becomes
$a_{1}(v_{1})+2a_{2}(v_{1})\cdot x=x+v_{1},$
and thus $a_{2}(v_{1})=\frac{1}{2}$ and $a_{1}(v_{1})=v_{1}$. Then Theorem 3.2 gives:
$Z_{(1)}(s;v_{1};\bm{p})=\exp\Big{[}s\cdot\Big{(}v_{1}P_{-1}^{(1)}+\frac{1}{2}P_{-1}^{(2)}\Big{)}\Big{]}\cdot 1,$
where $t_{n}=\frac{p_{n}}{n}$ and
$P_{-1}^{(1)}=t_{1},\qquad P_{-1}^{(2)}=\sum_{n\geq 1}2(n+1)t_{n+1}\frac{\partial}{\partial t_{n}}.$
Example 3.2.
Let $r=2$. In this special case, the above theorem recovers the cut-and-join representation for the dessin partition function $Z_{\text{dessin}}(s;v_{1},v_{2};\bm{p})$ derived by Kazarian and Zograf in [24].
In fact, the polynomials $a_{k}(\bm{v})=a_{k}(v_{1},v_{2})$ (where $k=1,2,3$) are determined by:
$a_{1}(\bm{v})+2a_{2}(\bm{v})\cdot x+3a_{3}(\bm{v})\cdot x(x-1)=(x+v_{1})(x+v_{2}),$
and thus:
$a_{1}(\bm{v})=v_{1}v_{2},\qquad a_{2}(\bm{v})=\frac{v_{1}+v_{2}+1}{2},\qquad a_{3}(\bm{v})=\frac{1}{3}.$
Then Theorem 3.2 gives a cut-and-join type representation for $Z_{\text{dessin}}(s;v_{1},v_{2};\bm{p})$:
(39) $Z_{\text{dessin}}(s;v_{1},v_{2};\bm{p})=\exp\Big{[}s\cdot\Big{(}v_{1}v_{2}P_{-1}^{(1)}+\frac{v_{1}+v_{2}+1}{2}P_{-1}^{(2)}+\frac{1}{3}P_{-1}^{(3)}\Big{)}\Big{]}\cdot 1,$
where $t_{n}=\frac{p_{n}}{n}$ and
$P_{-1}^{(3)}=3\sum_{k,l\geq 1}\Big{(}klt_{k}t_{l}\frac{\partial}{\partial t_{k+l-1}}+(k+l+1)t_{k+l+1}\frac{\partial}{\partial t_{k}}\frac{\partial}{\partial t_{l}}\Big{)}-3\sum_{k\geq 1}(k+1)t_{k+1}\frac{\partial}{\partial t_{k}}.$
Therefore the cut-and-join operator $\sum_{k=1}^{3}a_{k}(v_{1},v_{2})P_{-1}^{(k)}$ is:
(40) $\begin{split}\sum_{k=1}^{3}a_{k}(\bm{v})P_{-1}^{(k)}=&\sum_{k,l\geq 1}\Big{(}klt_{k}t_{l}\frac{\partial}{\partial t_{k+l-1}}+(k+l+1)t_{k+l+1}\frac{\partial}{\partial t_{k}}\frac{\partial}{\partial t_{l}}\Big{)}\\ &+(v_{1}+v_{2})\sum_{n\geq 1}(n+1)t_{n+1}\frac{\partial}{\partial t_{n}}+v_{1}v_{2}\cdot t_{1}.\end{split}$
This is the cut-and-join operator in [24, Corollary 1]. One can also use the notations of the free bosons to simplify the expressions. Let $\{\alpha_{n}\}_{n\in{\mathbb{Z}}}$ be the bosonic operators on ${\mathbb{C}}[[\bm{t}]]$:
(41) $\alpha_{n}=\begin{cases}(-n)t_{-n},\qquad&n<0;\\ 0,&n=0;\\ \frac{\partial}{\partial t_{n}},&n>0,\end{cases}$
then:
$\sum_{k=1}^{3}a_{k}(\bm{v})P_{-1}^{(k)}=\frac{1}{3}\sum_{i+j+k=-1}:\alpha_{i}\alpha_{j}\alpha_{k}:+\frac{v_{1}+v_{2}}{2}\sum_{i+j=-1}:\alpha_{i}\alpha_{j}:+v_{1}v_{2}\alpha_{-1}.$
where ::::$\quad$:: : is the normal ordering of bosonic operators.
Example 3.3.
Let $r=3$. In this case, the polynomials $a_{k}(v_{1},v_{2},v_{3})$ (where $1\leq k\leq 4$) are determined by:
$a_{1}(\bm{v})+2a_{2}(\bm{v})\cdot x+3a_{3}(\bm{v})\cdot x(x-1)+4a_{4}(\bm{v})\cdot x(x-1)(x-2)=\prod_{i=1}^{3}(x+v_{i}),$
and thus:
$\begin{split}a_{1}(\bm{v})=e_{3}(\bm{v}),\quad a_{2}(\bm{v})=\frac{e_{2}(\bm{v})+e_{1}(\bm{v})+1}{2}\quad a_{3}(\bm{v})=\frac{e_{1}(\bm{v})+3}{3},\quad a_{4}(\bm{v})=\frac{1}{4},\end{split}$
where $e_{i}(\bm{v})=e_{i}(v_{1},v_{2},v_{3})$ is the elementary symmetric polynomial of degree $i$ in $v_{1},v_{2},v_{3}$. Then Theorem 3.2 gives a cut-and-join type representation:
$\begin{split}&Z_{(3)}(s;v_{1},v_{2},v_{3};\bm{p})\\ =&\exp\Big{[}s\cdot\Big{(}e_{3}(\bm{v})P_{-1}^{(1)}+\frac{e_{2}(\bm{v})+e_{1}(\bm{v})+1}{2}P_{-1}^{(2)}+\frac{e_{1}(\bm{v})+3}{3}P_{-1}^{(3)}+\frac{1}{4}P_{-1}^{(4)}\Big{)}\Big{]}\cdot 1,\end{split}$
where $t_{n}=\frac{p_{n}}{n}$. One can also represent the operators $P_{-1}^{(k)}$ and $\sum_{k=1}^{4}a_{k}(\bm{v})P_{-1}^{(k)}$ in terms of the bosonic operators $\{\alpha_{n}\}$. Here we omit the details and write down the result (cf. [2, §8.4]):
$\begin{split}Z_{(3)}(s;\bm{v};\bm{p})=\exp\Big{[}s\Big{(}&e_{3}(\bm{v})\alpha_{-1}+\frac{e_{2}(\bm{v})}{2}\sum_{i+j=-1}:\alpha_{i}\alpha_{j}:+\frac{e_{1}(\bm{v})}{3}\sum_{i+j+k=-1}:\alpha_{i}\alpha_{j}\alpha_{k}:\\ &+\frac{1}{4}\sum_{i+j+k+l=-1}:\alpha_{i}\alpha_{j}\alpha_{k}\alpha_{l}:+\sum_{l\geq 1}\frac{l(l+1)}{2}\alpha_{-l-1}\alpha_{l}\Big{)}\Big{]}\cdot 1.\end{split}$
Example 3.4.
Here we omit the details and write down the cut-and-join type representations for $r=4$ and $r=5$:
$\begin{split}&Z_{(4)}(s;\bm{v};\bm{p})=\exp\Big{[}s\Big{(}e_{4}(\bm{v})P_{-1}^{(1)}+\frac{1}{2}\big{(}e_{0}(\bm{v})+e_{1}(\bm{v})+e_{2}(\bm{v})+e_{3}(\bm{v})\big{)}P_{-1}^{(2)}\\ &\quad+\frac{1}{3}\big{(}7e_{0}(\bm{v})+3e_{1}(\bm{v})+e_{2}(\bm{v})\big{)}P_{-1}^{(3)}+\frac{1}{4}\big{(}6e_{0}(\bm{v})+e_{1}(\bm{v})\big{)}P_{-1}^{(4)}+\frac{e_{0}(\bm{v})}{5}P_{-1}^{(5)}\Big{)}\Big{]}\cdot 1,\end{split}$
and:
$\begin{split}&Z_{(5)}(s;\bm{v};\bm{p})=\exp\Big{[}s\Big{(}e_{5}(\bm{v})P_{-1}^{(1)}+\frac{1}{2}\big{(}e_{0}(\bm{v})+e_{1}(\bm{v})+e_{2}(\bm{v})+e_{3}(\bm{v})+e_{4}(\bm{v})\big{)}P_{-1}^{(2)}\\ &\quad+\frac{1}{3}\big{(}15e_{0}(\bm{v})+7e_{1}(\bm{v})+3e_{2}(\bm{v})+e_{3}(\bm{v})\big{)}P_{-1}^{(3)}+\frac{1}{4}\big{(}25e_{0}(\bm{v})+6e_{1}(\bm{v})+\\ &\quad e_{2}(\bm{v})\big{)}P_{-1}^{(4)}+\frac{1}{5}\big{(}10e_{0}(\bm{v})+e_{1}(\bm{v})\big{)}P_{-1}^{(5)}+\frac{1}{6}e_{0}(\bm{v})P_{-1}^{(6)}\Big{)}\Big{]}\cdot 1,\end{split}$
where we use the convention $e_{0}(\bm{v})=1$.
3.4. Explicit formula for $a_{k}({\bf v})$
From the above examples one immediately guesses that the functions $a_{k}(\bm{v})=a_{k}(v_{1},\cdots,v_{r})$ are all linear combinations of the elementary symmetric functions $e_{i}(\bm{v})$ where $0\leq i\leq r$. In this subsection, we give an explicit formula for these $a_{k}(\bm{v})$.
Proposition 3.2.
Let $r\geq 1$. Then the functions $\{a_{k}(\bm{v})\}_{k=1}^{r}$ are given by:
(42) $a_{k}(\bm{v})=\frac{(-1)^{k-1}}{k!}e_{r}(\bm{v})+\frac{1}{k}\sum_{j=0}^{r}\Big{(}\sum_{i=1}^{k-1}\frac{(-1)^{i+k-1}\cdot i^{r-j}}{i!\cdot(k-1-i)!}\Big{)}e_{j}(\bm{v}),$
where $e_{j}(\bm{v})$ is the elementary symmetric polynomial of degree $j$ in $v_{1},v_{2},\cdots,v_{r}$ for $1\leq j\leq r$, and $e_{0}(\bm{v})=1$.
Proof.
In this proof, we need to remember the integer $r$ we choose. We shall denote $a_{k}(\bm{v})$ for $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$ by $a_{(r),k}(\bm{v})$. Recall that $a_{(r),k}(\bm{v})$ is determined by:
(43) $\sum_{k=1}^{r+1}ka_{(r),k}(\bm{v})\cdot\prod_{j=0}^{k-2}(x-j)=\prod_{i=1}^{r}(x+v_{i})=\sum_{j=0}^{r}x^{r-j}\cdot e_{j}(\bm{v}).$
Take $x=0$ in this equality, we obtain:
(44) $a_{(r),1}(\bm{v})=v_{1}v_{2}\cdots v_{r}=e_{r}(\bm{v}),$
for every $r$. Then we take $x=1$ and obtain:
(45) $\begin{split}a_{(r),2}(\bm{v})=&\frac{1}{2}\Big{(}(1+v_{1})(1+v_{2})\cdots(1+v_{r})-a_{(r),1}(\bm{v})\Big{)}\\ =&\frac{1}{2}\big{(}e_{0}(\bm{v})+e_{1}(\bm{v})+\cdots+e_{r-1}(\bm{v})\big{)}.\end{split}$
And Similarly, taking $x=2$ gives:
(46) $\begin{split}a_{(r),3}(\bm{v})=&\frac{1}{3}\sum_{i=0}^{r-1}(2^{r-i-1}-1)e_{i}(\bm{v})\\ =&\frac{1}{3}\Big{(}e_{r-2}(\bm{v})+3e_{r-3}(\bm{v})+7e_{r-4}(\bm{v})+\cdots+(2^{r-1}-1)e_{0}(\bm{v})\Big{)}.\end{split}$
Repeat this procedure by taking $x=3,4,5,\cdots$ successively, and inductively we may see that:
  • 1)
    Every $a_{(r),k}(\bm{v})$ is a linear combination of $\{e_{i}(\bm{v})\}_{i=0}^{r}$.
  • 2)
    Fix two integers $j\geq 0$ and $k\geq 1$. When $r>\max\{j,k\}$, the coefficient of $e_{r-j}(\bm{v})$ in the expansion of $a_{(r),k}$ is independent of $r$. More precisely, if $a_{(r),k}(\bm{v})$ is of the form:
    $a_{(r),k}(\bm{v})=C_{r}e_{r}(\bm{v})+C_{r-1}e_{r-1}(\bm{v})+\cdots+C_{0}e_{0}(\bm{v}),$
    then $a_{(r+1),k}(\bm{v})$ must be of the form:
    $a_{(r),k}(\bm{v})=C_{r}e_{r+1}(\bm{v})+C_{r-1}e_{r}(\bm{v})+\cdots+C_{0}e_{1}(\bm{v})+Ce_{0}(\bm{v})$
    for some constant $C$.
Thus in order to prove the proposition, we only need to show that the coefficient of $e_{0}(\bm{v})=1$ in $a_{(r),k}(\bm{v})$ is:
(47) $\frac{1}{k}\sum_{i=1}^{k-1}\frac{(-1)^{i+k-1}\cdot i^{r}}{i!\cdot(k-1-i)!}.$
We prove this by induction on $k$. The cases $k=1,2,3$ can be checked directly using (44)-(46). Now assume this holds for $k=1,2,\cdots,n$ (where $n\leq r$), and consider the coefficient of $e_{0}(\bm{v})$ in $a_{(r),n+1}(\bm{v})$. We denote this coefficient by $A$. Now take $x=n$ in (43) and compute the coefficients of $e_{0}(\bm{v})$, and we obtain:
$\sum_{k=1}^{n}\Big{(}\sum_{i=1}^{k-1}\frac{(-1)^{i+k-1}\cdot i^{r}}{i!\cdot(k-1-i)!}\Big{)}\Big{(}\prod_{j=0}^{k-2}(n-j)\Big{)}+(n+1)!\cdot A=n^{r}.$
Notice that:
$\begin{split}&\sum_{k=1}^{n}\Big{(}\sum_{i=1}^{k-1}\frac{(-1)^{i+k-1}\cdot i^{r}}{i!\cdot(k-1-i)!}\Big{)}\Big{(}\prod_{j=0}^{k-2}(n-j)\Big{)}\\ =&\sum_{k=1}^{n}\sum_{i=1}^{k-1}\frac{(-1)^{i+k-1}\cdot i^{r}\cdot n!}{i!\cdot(k-1-i)!\cdot(n-k+1)!}\\ =&\sum_{i=1}^{n-1}\frac{(-1)^{i-1}\cdot i^{r}\cdot n!}{i!}\sum_{k=i+1}^{n}\frac{(-1)^{k}}{(k-1-i)!\cdot(n-k+1)!},\end{split}$
where
$\sum_{k=i+1}^{n}\frac{(-1)^{k}}{(k-1-i)!\cdot(n-k+1)!}=-\frac{(-1)^{n+1}}{(n-i)!\cdot 0!}=\frac{(-1)^{n}}{(n-i)!},$
and therefore we have:
$A=\frac{1}{(n+1)!}\Big{(}n^{r}-\sum_{i=1}^{n-1}\frac{(-1)^{i-1}\cdot i^{r}\cdot n!}{i!}\cdot\frac{(-1)^{n}}{(n-i)!}\Big{)}=\frac{1}{n+1}\sum_{i=1}^{n}\frac{(-1)^{i+n}\cdot i^{r}}{i!\cdot(n-i)!},$
which completes the proof. ∎
3.5. KP integrability and affine coordinates
It is known in literatures that the operators $P_{m}^{(k)}$ belong to the algebra $W_{1+\infty}$ which gives the symmetry of the KP hierarchy, see e.g. [14, 27]. Therefore for every $r\geq 1$, the partition function
$Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})=e^{s\cdot\sum_{k=1}^{r+1}a_{k}(\bm{v})P_{-1}^{(k)}}(1)$
is a tau-function of the KP hierarchy with time variables $\bm{t}=(t_{1},t_{2},t_{3},\cdots)$, where $t_{n}=\frac{p_{n}}{n}$. In this subsection, we write down the affine coordinates of this tau-function on the Sato Grassmannian.
First we recall some basics of the affine coordinates. In Sato’s theory [7, 32], the space of all formal power series tau-functions of the KP hierarchy is a semi-infinite dimensional Grassmannian called the Sato Grassmannian. Let $\tau(\bm{t})$ be a tau-function of the KP hierarchy with time variables $\bm{t}=(t_{1},t_{2},\cdots)$ satisfying the initial condition $\tau(\bm{0})=1$, then it lies on the big cell of the Sato Grassmannian. The affine coordinates $\{a_{n,m}\in{\mathbb{C}}\}_{n,m\geq 0}$ are natural coordinates on the big cell, and thus a tau-function $\tau(\bm{t})$ satisfying $\tau(\bm{0})=1$ can be completely described in terms of its affine coordinates, see e.g. [7, 35, 4] for details. In particular, the tau-function admits an expansion of Schur functions whose coefficients are determinants of its affine coordinates $\{a_{n,m}\}_{n,m\geq 0}$:
(48) $\tau(\bm{t})=\sum_{\mu\in{\mathcal{P}}}(-1)^{n_{1}+\cdots+n_{k}}\cdot\det\big{(}a_{n_{i},m_{j}}\big{)}_{1\leq i,j\leq k}\cdot s_{\mu}(\bm{t}),$
where $\mu=(m_{1},\cdots,m_{k}|n_{1},\cdots,n_{k})$ is the Frobenius notation of the partition $\mu$. In particular, $a_{n,m}$ is $(-1)^{n}$ times the coefficient of $s_{(m|n)}=s_{(m+1,1^{n})}$.
Now by taking the coefficient of $s_{(m|n)}$ in the Schur function expansion (29) for $Z_{(r)}(s;\bm{v};\bm{p})$, we see that:
Proposition 3.3.
Let $r\geq 1$. The affine coordinates $\{a_{n,m}^{(r)}\}_{n,m\geq 0}$ of the tau-function $Z_{(r)}(s;\bm{v};\bm{p})$ are:
(49) $\begin{split}a_{n,m}^{(r)}=(-1)^{n}\cdot\frac{s^{m+n+1}}{(m+n+1)\cdot m!\cdot n!}\cdot\prod_{i=1}^{r}\prod_{j=-n}^{m}(v_{i}+j).\end{split}$
Then one can use Zhou’s formula (see [35, Theorem 5.3]) to compute the connected $n$-point functions of the tau-function $Z_{(r)}(s;\bm{v};\bm{p})$:
Theorem 3.3 ([35]).
Let $\tau(\bm{t})$ be a tau-function of the KP hierarchy satisfying $\tau(\bm{0})=1$, and let $\{a_{n,m}\}_{n,m\geq 0}$ be its affine coordinates. Denote:
(50) $A(z,w)=\sum_{n,m\geq 0}a_{n,m}z^{-n-1}w^{-m-1},$
then the connected $n$-point functions are given by:
(51) $\begin{split}&\sum_{j_{1},\cdots,j_{n}\geq 0}\frac{\partial^{n}\log\tau(\bm{t})}{\partial t_{j_{1}}\cdots\partial t_{j_{n}}}\Big{|}_{\bm{t}=0}\cdot z_{1}^{-j_{1}-1}\cdots z_{n}^{-j_{n}-1}\\ =&(-1)^{n-1}\cdot\sum_{\text{$n$-cycles }\sigma}\Big{(}\prod_{i=1}^{n}\widehat{A}(z_{\sigma(i),\sigma(i+1)})\Big{)}-\frac{\delta_{n,2}}{(z_{1}-z_{2})^{2}},\end{split}$
where $\sigma(n+1)=\sigma(n)$, and
(52) $\widehat{A}(z_{i},z_{j})=\begin{cases}i_{z_{i},z_{j}}\frac{1}{z_{i}-z_{j}}+A(z_{i},z_{j}),\qquad&i<j;\\ A(z_{i},z_{i}),&i=j;\\ i_{z_{j},z_{i}}\frac{1}{z_{i}-z_{j}}+A(z_{i},z_{j}),&i>j,\end{cases}$
and $i_{z,w}\frac{1}{z+w}=\sum_{k\geq 0}(-1)^{k}z^{-1-k}w^{k}$.
In our case, the free energy $\log Z_{(r)}(s;\bm{v};\bm{p})$ is:
$\begin{split}&\log Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})\\ =&\sum_{k_{1},\cdots,k_{r}=1}^{\infty}\sum_{\mu\in{\mathcal{P}}}\Big{(}\prod_{i=1}^{r}v_{i}^{k_{i}}\Big{)}s^{|\mu|}N_{k_{1},\cdots,k_{r}}^{\circ}(\mu)\cdot\mu_{1}t_{\mu_{1}}\cdots\mu_{l(\mu)}t_{\mu_{l(\mu)}}.\end{split}$
For simplicity we may take $s=1$, and then we obtain:
Proposition 3.4.
Let $\mu=(\mu_{1},\cdots,\mu_{l})$ be a partition, then:
(53) $\begin{split}&\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}(\mu)\cdot\prod_{i=1}^{r}v_{i}^{k_{i}}\\ =&\frac{1}{z_{\mu}}\cdot\Big{(}(-1)^{n-1}\sum_{\text{$n$-cycles }\sigma}\prod_{i=1}^{n}\widehat{A}_{(r)}(z_{\sigma(i),\sigma(i+1)})-\frac{\delta_{n,2}}{(z_{1}-z_{2})^{2}}\Big{)}_{[z^{-\mu-1}]},\end{split}$
where
$z_{\mu}=\prod_{i\geq 1}i^{m_{i}(\mu)}\cdot m_{i}(\mu)!,$
and $(\quad)_{[z^{-\mu-1}]}$ means taking the coefficient of $z_{1}^{-\mu_{1}-1}\cdots z_{l}^{-\mu_{l}-1}$. Here
(54) $A_{(r)}(z,w)=\sum_{n,m\geq 0}\frac{(-1)^{n}}{(m+n+1)\cdot m!\cdot n!}\Big{(}\prod_{i=1}^{r}\prod_{j=-n}^{m}(v_{i}+j)\Big{)}z^{-n-1}w^{-m-1},$
and the series $\widehat{A}_{(r)}(z,w)$ is constructed from $A_{(r)}(z,w)$ by applying (52).
Example 3.5.
First consider the connected $1$-point function. Take $\mu=(n)$, then the above proposition gives:
(55) $\begin{split}\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(n)\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}=&\frac{1}{n}\cdot\Big{(}A_{(r)}(z,z)\Big{)}_{[z^{-n-1}]}\\ =&\frac{1}{n}\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\cdot\prod_{i=1}^{r}\prod_{j=-k}^{l}(v_{i}+j)}{(k+l+1)\cdot k!\cdot l!}.\end{split}$
In particular,
$\begin{split}\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(1)\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}=&v_{1}v_{2}\cdots v_{r}=e_{r}(\bm{v}),\\ \sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(2)\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}=&\frac{1}{4}\Big{(}-\prod_{i=1}^{r}(v_{i}-1)v_{i}+\prod_{i=1}^{r}v_{i}(v_{i}+1)\Big{)}\\ =&\frac{1}{4}e_{r}(\bm{v})\cdot\sum_{i=0}^{r}\Big{(}1-(-1)^{r-i}\Big{)}e_{i}(\bm{v}),\end{split}$
and for $n=3$,
$\begin{split}&\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(3)\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}\\ =&\frac{1}{3}\Big{(}\frac{1}{6}\prod_{i=1}^{r}(v_{i}-2)(v_{i}-1)v_{i}-\frac{1}{3}\prod_{i=1}^{r}(v_{i}-1)v_{i}(v_{i}+1)+\frac{1}{6}\prod_{i=1}^{r}v_{i}(v_{i}+1)(v_{i}+2)\Big{)}\\ =&\frac{e_{r}(\bm{v})}{18}\cdot\sum_{i,j=0}^{r}\Big{(}\big{(}(-1)^{i+j}+1\big{)}2^{r-j}-2\cdot(-1)^{r-i}\Big{)}e_{i}(\bm{v})e_{j}(\bm{v}).\end{split}$
Example 3.6.
Now consider the $2$-point function. Take $\mu=(n_{1},n_{2})$ where $n_{1}\geq n_{2}\geq 1$, then:
(56) $\begin{split}&\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(n_{1},n_{2})\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}\\ =&\frac{1}{n_{1}n_{2}(1+\delta_{n_{1},n_{2}})}\Big{(}-\widehat{A}_{(r)}(z_{1},z_{2})\widehat{A}_{(r)}(z_{2},z_{1})-\frac{1}{(z_{1}-z_{2})^{2}}\Big{)}_{[z_{1}^{-n_{1}-1}z_{2}^{-n_{2}-2}]},\end{split}$
where
$\begin{split}&-\widehat{A}_{(r)}(z_{1},z_{2})\widehat{A}_{(r)}(z_{2},z_{1})-\frac{1}{(z_{1}-z_{2})^{2}}\\ =&i_{z_{1},z_{2}}\frac{1}{z_{1}-z_{2}}\cdot\Big{(}A_{(r)}(z_{1},z_{2})-A_{(r)}(z_{2},z_{1})\Big{)}-A_{(r)}(z_{1},z_{2})A_{(r)}(z_{2},z_{1}),\end{split}$
and then:
$\begin{split}&\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(n_{1},n_{2})\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}\\ =&\frac{1}{n_{1}n_{2}(1+\delta_{n_{1},n_{2}})}\Big{[}\sum_{k=0}^{n_{1}-1}\frac{(-1)^{n_{1}-1-k}}{(n_{1}+n_{2})\cdot(n_{1}-1-k)!\cdot(n_{2}+k)!}\prod_{i=1}^{r}\prod_{j=k+1-n_{1}}^{n_{2}+k}(v_{i}+j)\\ &\qquad\quad-\sum_{k=0}^{n_{1}-1}\frac{(-1)^{n_{2}+k}}{(n_{1}+n_{2})\cdot(n_{1}-1-k)!\cdot(n_{2}+k)!}\prod_{i=1}^{r}\prod_{j=-n_{2}-k}^{n_{1}-1-k}(v_{i}+j)\\ &\qquad\quad-\sum_{l=0}^{n_{1}-1}\sum_{k=0}^{n_{2}-1}\Big{(}\frac{(-1)^{n_{1}-1-l}\cdot\prod_{i=1}^{r}\prod_{j=l+1-n_{1}}^{n_{2}-1-k}(v_{i}+j)}{(n_{1}+n_{2}-k-l-1)\cdot(n_{2}-1-k)!\cdot(n_{1}-1-l)!}\\ &\qquad\quad\cdot\frac{(-1)^{k}}{(l+k+1)\cdot l!\cdot k!}\cdot\prod_{i=1}^{r}\prod_{j=-k}^{l}(v_{i}+j)\Big{)}\Big{]}.\end{split}$
This formula will be useful in §5. Take $(n_{1},n_{2})=(1,1)$ and $(2,1)$, and we obtain:
$\begin{split}\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(1,1)\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}=\frac{1}{4}e_{r}(\bm{v})\cdot\sum_{i=0}^{r-1}\big{(}1+(-1)^{r-i}\big{)}e_{i}(\bm{v}),\end{split}$
and
$\begin{split}&\sum_{k_{1},\cdots,k_{r}=1}^{\infty}N_{k_{1},\cdots,k_{r}}^{\circ}\Big{(}(2,1)\Big{)}\prod_{i=1}^{r}v_{i}^{k_{i}}\\ =&\frac{e_{r}(\bm{v})}{12}\sum_{i,j=0}^{r}\big{(}1-(-1)^{i+j}\big{)}\cdot 2^{r-j}e_{i}(\bm{v})e_{j}(\bm{v})+\frac{e_{r}(\bm{v})^{2}}{4}\sum_{i=0}^{r}\big{(}(-1)^{r-i}-1\big{)}e_{i}(\bm{v}).\end{split}$
4. A Polynomial Structure for the Disconnected Correlators
In this section, we prove a polynomial structure for the disconnected correlators of the generalized dessin partition function $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$. This polynomial structure is inspired by a result of Stanley [33] (see also [20, 31] for some generalizations). Now fix a positive integer $r$, and our main result in this section is:
Theorem 4.1.
Fix two partitions of integers $\lambda=(\lambda_{1},\cdots,\lambda_{k})$ (where $k\leq r$) and $\mu=(\mu_{1},\cdots,\mu_{l})$. Then there is a polynomial $A_{(r)}^{\lambda,\mu}(x)\in{\mathbb{Q}}[x]$, such that the disconnected correlators of $Z_{(r)}(s;v_{1},\cdots,v_{r};\bm{p})$ satisfies:
(57) $n!\cdot N_{n-\lambda_{1},\cdots,n-\lambda_{k},n,\cdots,n}^{\bullet}\big{(}(\mu_{1},\cdots,\mu_{l},1^{n-|\mu|})\big{)}=A_{(r)}^{\lambda,\mu}(n)$
for every positive integer $n$ with $n\geq|\mu|$.
Remark 4.1.
Theorem 4.1 is actually a simple modification of Stanley’s result [33, Theorem 2.1] which gives the special case $\mu=(\emptyset)$.
Remark 4.2.
For simplicity, in what follows we will say that the left-hand side of (57) is a polynomial in $n$ even though here $n$ only takes values in positive integers greater than $|\mu|$. When we say $f(n)$ is a polynomial in a positive integer $n$, we always mean there exists a polynomial $g(x)$ such that $g(n)=f(n)$ for every $n$.
Proof.
We use Stanley’s method to prove Theorem 4.1. Recall that there is a group-theoretic description (13) of the Hurwitz numbers. From this description we know that:
(58) $\begin{split}&n!\cdot N_{n-\lambda_{1},\cdots,n-\lambda_{k},n,\cdots,n}^{\bullet}\big{(}(\mu_{1},\cdots,\mu_{l},1^{n-|\mu|})\big{)}\\ =&\Bigg{\{}(\alpha_{1},\cdots,\alpha_{k},\alpha)\in(S_{n})^{k+1}\Bigg{|}\begin{array}[]{l}\alpha_{1}\cdots\alpha_{k}\cdot\alpha=e\in S_{n};\\ \text{$\alpha_{i}$ has $n-\lambda_{i}$ cycles;}\\ \text{$\alpha$ is of type $(\mu_{1},\cdots,\mu_{l},1^{n-|\mu|})$}\end{array}\Bigg{\}}.\end{split}$
Here $\alpha_{1},\cdots,\alpha_{k},\alpha\in S_{n}$ are bijections from the set of indices $[n]=\{1,2,\cdots,n\}$ to itself. Now denote:
$\begin{split}&F_{i}=\big{\{}j\in[n]\big{|}\text{$j$ is not fixed by $\alpha_{i}$}\big{\}},\qquad 1\leq i\leq k;\\ &F_{0}=\big{\{}j\in[n]\big{|}\text{$j$ is not fixed by $\alpha$}\big{\}}.\end{split}$
Then $\alpha_{i}:[n]\to[n]$ has $n-|F_{i}|$ fixed points, i.e., $n-|F_{i}|$ cycles of length one, and thus has $(n-\lambda_{i})-(n-|F_{i}|)$ cycles of length at least two. Therefore
$n\geq 2\cdot\big{(}(n-\lambda_{i})-(n-|F_{i}|)\big{)}+1\cdot(n-|F_{i}|),$
or equivalently, $|F_{i}|\leq 2\lambda_{i}$ for every $1\leq i\leq k$. Moreover, $\alpha$ has at least $n-|\mu|$ cycles of length one, and thus $n-|F_{0}|\geq n-|\mu|$, or equivalently, $|F_{0}|\leq|\mu|$. Denote:
$F=F_{0}\cup F_{1}\cup\cdots\cup F_{k}\subset[n],$
then one has:
$|F|\leq\sum_{i=0}^{k}|F_{i}|\leq|\mu|+2|\lambda|.$
In conclusion, the size of $F$ has the upper bound $|\mu|+2|\lambda|$, which is independent of the choice of the $d$ numbers in $[n]$.
Denote $d=|F|$. For a fixed $d$, there are $\binom{n}{d}$ choices of such a subset $F\subset[n]$; and for a fixed choice of $F\subset[n]$, the $(n-d)$ elements in $[n]\backslash F$ are fixed by all $\alpha_{i}$ and $\alpha$. Hence the bijections $\alpha_{1},\cdots,\alpha_{k},\alpha$ on $[n]$ induce bijections $\tilde{\alpha}_{1},\cdots,\tilde{\alpha}_{k},\tilde{\alpha}$ on the set $F$, and here $\tilde{\alpha}_{i}$ has $(n-\lambda_{i})-(n-d)=(d-\lambda_{i})$ cycles. Denote:
$\begin{split}T_{d}=&\Bigg{\{}(\tilde{\alpha}_{1},\cdots,\tilde{\alpha}_{k},\tilde{\alpha})\in(S_{d})^{k+1}\Bigg{|}\begin{array}[]{l}\tilde{\alpha}_{1}\cdots\tilde{\alpha}_{k}\cdot\tilde{\alpha}=\operatorname{id}_{[d]};\quad\text{$\alpha_{i}$ has $d-\lambda_{i}$ cycles;}\\ \text{$\alpha$ is of type $(\mu_{1},\cdots,\mu_{l},1^{d-|\mu|})$;}\\ \text{no element in $[d]$ is fixed by all $\alpha_{i}$ and $\alpha$}\end{array}\Bigg{\}}.\end{split}$
Clearly $|T_{d}|$ is independent of $n$, and then by (58) we see that when $n\geq|\mu|+2|\lambda|$,
(59) $n!\cdot N_{n-\lambda_{1},\cdots,n-\lambda_{k},n,\cdots,n}^{\bullet}\big{(}(\mu_{1},\cdots,\mu_{l},1^{n-|\mu|})\big{)}=\sum_{d=0}^{|\mu|+2|\lambda|}|T_{d}|\cdot\binom{n}{d}$
is a polynomial in $n$ whose coefficients are all rational numbers. For positive integers $n$ with $n<|\mu|+2|\lambda|$, the above relation still holds where by definition $\binom{n}{d}=0$ for $d>n$. Thus the conclusion holds. ∎
5. A Conjectural Polynomial Structure for the Connected Correlators
In this section we propose a conjectural polynomial structure for the connected correlators $N_{k_{1},\cdots,k_{r}}^{\circ}(\mu)$ and prove some special cases. This conjecture is inspired by the polynomial structure for various types of Hurwitz numbers in literature, see e.g. [13, 9, 12, 17, 18, 25]. Actually, this polynomiality is a generalization of many known cases. For examples, the $k_{1}=\cdots k_{r}=1$ case is the one-part double Hurwitz numbers studied in [18], and the $r=1$ case gives the strictly monotone Hurwitz numbers studied in [9, 25].
Conjecture 5.1.
Fix some positive integers $l$ and $k_{1},k_{2},\cdots,k_{r}$. There exists a polynomial $A_{k_{1},\cdots,k_{r}}(x_{1},\cdots x_{l})\in{\mathbb{Q}}[x_{1},\cdots,x_{l}]$ such that
(60) $A_{k_{1},\cdots,k_{r}}(\mu_{1},\cdots\mu_{l})=z_{\mu}\cdot N_{|\mu|-k_{1},|\mu|-k_{2},\cdots,|\mu|-k_{r-1},k_{r}}^{\circ}(\mu)$
for every partition $\mu=(\mu_{1},\cdots,\mu_{l})$ satisfying $|\mu|>\max\{k_{1},\cdots,k_{r}\}$.
In the following subsections, we will give some evidences by numerical computations and prove the one-point case of this conjecture.
5.1. Examples of the $1$-point and $2$-point correlators
In this subsection, we give some examples of the connected correlators.
First consider the one-point (i.e., $l(\mu)=1$) case. Let $\mu=(n)$, and we need to show that $n\cdot N_{n-k_{1},\cdots,n-k_{r-1},k_{r}}^{\circ}\big{(}(n)\big{)}$ is a polynomial in $n$ (when $n>k_{i}$ for $i=1,\cdots,r-1$ and $n\geq k_{r}$). By the results of Example 3.5, we know that
(61) $\begin{split}&n\cdot N_{n-k_{1},\cdots,n-k_{r-1},k_{r}}^{\circ}\big{(}(n)\big{)}\\ =&\Big{(}\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\prod_{i=1}^{r}\prod_{j=-k}^{l}(v_{i}+j)}{(k+l+1)\cdot k!\cdot l!}\Big{)}_{[v_{1}^{n-k_{1}}\cdots v_{r-1}^{n-k_{r-1}}v_{r}^{k_{r}}]},\end{split}$
where $(\quad)_{[v_{1}^{n-k_{1}}\cdots v_{r-1}^{n-k_{r-1}}v_{r}^{k_{r}}]}$ means taking the coefficient of $v_{1}^{n-k_{1}}\cdots v_{r-1}^{n-k_{r-1}}v_{r}^{k_{r}}$.
Example 5.1.
First consider the simplest case $r=1$. In this case the connected correlator (60) is simply $n\cdot N_{k}^{\circ}\big{(}(n)\big{)}$ for a fixed $k$. From (61) one easily checks that:
$\begin{split}&nN_{1}^{\circ}\big{(}(n)\big{)}=\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}}{n\cdot k!\cdot l!}\Big{(}\prod_{j=-k}^{-1}j\Big{)}\Big{(}\prod_{j=1}^{l}j\Big{)}=1,\\ &nN_{2}^{\circ}\big{(}(n)\big{)}=\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\cdot(-1)^{k}\cdot k!\cdot l!}{n\cdot k!\cdot l!}\Big{(}\sum_{j=-k}^{-1}\frac{1}{j}+\sum_{j=1}^{l}\frac{1}{j}\Big{)}=0.\end{split}$
In general, we have:
(62) $N_{k}^{\circ}\big{(}(n)\big{)}=0,\qquad\forall k>1.$
This can be proved by the Riemann Hurwitz formula (12). In fact, by (12) one has:
$2g-2=n\cdot(2-2)-(k+1),$
and thus $12g+k=$ 1. Here $g$ is the genus of a connected Riemann surface and hence $g\geq 0$, and thus there is no such branched covering when $k>1$.
Example 5.2.
Now we consider some examples for $r=2$. We have:
(63) $\begin{split}&nN_{n-1,1}^{\circ}\big{(}(n)\big{)}=0,\\ &nN_{n-1,2}^{\circ}\big{(}(n)\big{)}=\frac{n(n-1)}{2}=\binom{n}{2},\\ &nN_{n-2,1}^{\circ}\big{(}(n)\big{)}=\frac{1}{24}(n-2)(n-1)n(n+1)=\binom{n}{4}+\binom{n}{3}.\end{split}$
In fact, we take $k_{1}=k_{2}=1$ in (61) and get:
$\begin{split}nN_{n-1,1}^{\circ}\big{(}(n)\big{)}=&\Big{(}\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\prod_{j=-k}^{l}(v_{1}+j)(v_{2}+j)}{n\cdot k!\cdot l!}\Big{)}_{[v_{1}^{n-1}v_{2}]}\\ =&\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\cdot\big{(}\prod_{j=-k}^{-1}j\big{)}\big{(}\prod_{j=1}^{l}j\big{)}}{n\cdot k!\cdot l!}\Big{(}\sum_{j=-k}^{-1}j+\sum_{j=1}^{l}j\Big{)}=0.\end{split}$
And for $k_{1}=1$, $k_{2}=2$, we have:
$\begin{split}nN_{n-1,2}^{\circ}\big{(}(n)\big{)}=&\Big{(}\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\prod_{j=-k}^{l}(v_{1}+j)(v_{2}+j)}{n\cdot k!\cdot l!}\Big{)}_{[v_{1}^{n-1}v_{2}^{2}]}\\ =&\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\big{(}\prod_{j=-k}^{-1}j\big{)}\big{(}\prod_{j=1}^{l}j\big{)}}{n\cdot k!\cdot l!}\Big{(}\sum_{j=-k}^{-1}\frac{1}{j}+\sum_{j=1}^{l}\frac{1}{j}\Big{)}\Big{(}\sum_{j=-k}^{l}j\Big{)}\\ =&\frac{1}{n}\cdot\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\Big{(}\sum_{j=-k}^{l}j\Big{)}\Big{(}\sum_{j=-k}^{-1}\frac{1}{j}+\sum_{j=1}^{l}\frac{1}{j}\Big{)}.\end{split}$
Now let
$H_{k}=1+\frac{1}{2}+\frac{1}{3}+\cdots+\frac{1}{k}$
be the $k$-th harmonic number (and $H_{0}=0$), and then:
$nN_{n-1,2}^{\circ}\big{(}(n)\big{)}=\sum_{k=0}^{n-1}\Big{(}\frac{n-1}{2}-k\Big{)}\Big{(}H_{n-1-k}-H_{k}\Big{)}=\sum_{k=0}^{n-1}\Big{(}2k-n+1\Big{)}H_{k}.$
Now compute the difference $(n+1)N_{n,2}^{\circ}\big{(}(n+1)\big{)}-nN_{n-1,2}^{\circ}\big{(}(n)\big{)}$, we get:
$(n+1)N_{n,2}^{\circ}\big{(}(n+1)\big{)}-nN_{n-1,2}^{\circ}\big{(}(n)\big{)}=nH_{n}-(H_{n-1}+H_{n-2}+\cdots+H_{1}),$
where
$\sum_{i=1}^{n-1}H_{i}=\sum_{i=1}^{n-1}(n-i)\cdot\frac{1}{i}=nH_{n-1}-n+1,$
and thus:
$(n+1)N_{n,2}^{\circ}\big{(}(n+1)\big{)}-nN_{n-1,2}^{\circ}\big{(}(n)\big{)}=nH_{n}-(nH_{n-1}-n+1)=n.$
By this equality and the initial data $12N_{1,2}^{\circ}\big{(}(2)\big{)}=$ 1, one inductively proves the second equality in (63). The third equality in (63) can be proved similarly:
$\begin{split}nN_{n-2,1}^{\circ}\big{(}(n)\big{)}=&\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{1}{n}\Big{(}\sum_{1\leq i<j\leq k}ij+\sum_{1\leq i<j\leq l}ij-\sum_{i=1}^{k}\sum_{j=1}^{l}ij\Big{)}\\ =&\frac{1}{24}(n-2)(n-1)n(n+1).\end{split}$
Example 5.3.
Now we give some examples for $r=3$. First by the Riemann Hurwitz formula we know that $N_{n-k_{1},n-k_{2},k_{3}}^{\circ}\big{(}(n)\big{)}=0$ unless $k_{1}+k_{2}-k_{3}$ is an odd number greater than $-2$. In this case we have:
$\begin{split}nN_{n-1,n-1,1}^{\circ}\big{(}(n)\big{)}=&\frac{1}{12}n^{2}(n-1)(n+1)=2\binom{n}{4}+3\binom{n}{3}+\binom{n}{2},\\ nN_{n-1,n-2,2}^{\circ}\big{(}(n)\big{)}=&\frac{1}{24}(n-2)(n-1)^{2}n^{2}(n+1)\\ =&30\binom{n}{6}+60\binom{n}{5}+36\binom{n}{4}+6\binom{n}{3},\\ nN_{n-2,n-2,1}^{\circ}\big{(}(n)\big{)}=&\frac{1}{2880}(n-2)(n-1)^{2}n^{2}(n+1)(9n^{2}-n-18)\\ =&126\binom{n}{8}+392\binom{n}{7}+450\binom{n}{6}+230\binom{n}{5}+49\binom{n}{4}+3\binom{n}{3}.\end{split}$
Here we only prove the first equality since the proofs of other cases are all similar. Take $k_{1}=k_{2}=k_{3}=1$ in (61), and we see that:
$\begin{split}nN_{n-1,n-1,1}^{\circ}\big{(}(n)\big{)}=&\Big{(}\sum_{\begin{subarray}{c}k+l=n-1\\ k,l\geq 0\end{subarray}}\frac{(-1)^{k}\prod_{i=1}^{3}\prod_{j=-k}^{l}(v_{i}+j)}{(k+l+1)\cdot k!\cdot l!}\Big{)}_{[v_{1}^{n-1}v_{2}^{n-1}v_{3}]}\\ =&\frac{1}{n}\cdot\sum_{k=0}^{n-1}\Big{(}\sum_{j=-k}^{n-1-k}j\Big{)}^{2}\\ =&\frac{1}{12}n^{2}(n-1)(n+1).\end{split}$
Example 5.4.
Now we give some examples of the two-point connected correlators. Let $n_{1},n_{2}$ be two integers satisfying $n_{1}\geq n_{2}>0$. First consider the case $r=1$, we want to compute $N_{k}^{\circ}\big{(}(n_{1},n_{2})\big{)}$. By the Riemann-Hurwitz formula (12) one has:
$2g-2=(n_{1}+n_{2})\cdot(2-2)-(k+2),$
and thus $02g=-k<$ 0. Therefore there is no such branched covering, and we conclude that:
(64) $N_{k}^{\circ}\big{(}(n_{1},n_{2})\big{)}=0,\qquad\forall n_{1},n_{2}>0.$
Example 5.5.
Now consider the two-point connected correlators for $r=2$. The computations of
$z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-k_{1},k_{2}}^{\circ}\big{(}(n_{1},n_{2})\big{)}=n_{1}n_{2}(1+\delta_{n_{1},n_{2}})\cdot N_{n_{1}+n_{2}-k_{1},k_{2}}^{\circ}\big{(}(n_{1},n_{2})\big{)}$
is complicated, and it is not easy to derive explicit expressions for these correlators. In this case we can also apply the Riemann-Hurwitz formula and obtain:
$\begin{split}2g-2=&(n_{1}+n_{2})(3-2)-(n_{1}+n_{2}-k_{1}+k_{2}+2)\\ =&k_{1}-k_{2}-2,\end{split}$
and thus $N_{n_{1}+n_{2}-k_{1},k_{2}}^{\circ}\big{(}(n_{1},n_{2})\big{)}=0$ unless $k_{1}-k_{2}$ is a nonnegative even number. One can use the formulas in Example 3.6 to carry out some numerical data. The following conjectural formulas are checked for small $n_{1},n_{2}$:
$\begin{split}&z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-1,1}^{\circ}\big{(}(n_{1},n_{2})\big{)}=n_{1}\cdot n_{2};\\ &z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-2,2}^{\circ}\big{(}(n_{1},n_{2})\big{)}=\frac{1}{2}n_{1}n_{2}(n_{1}^{2}+n_{2}^{2}+n_{1}n_{2}-2n_{1}-2n_{2}+1);\\ &z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-3,1}^{\circ}\big{(}(n_{1},n_{2})\big{)}=\frac{n_{1}n_{2}}{24}\Big{(}n_{1}^{4}+(2n_{2}-4)n_{1}^{3}+(2n_{2}^{2}-6n_{2}+3)n_{1}^{2}\\ &\qquad\qquad+(2n_{2}^{3}-6n_{2}^{2}+2n_{2}+4)n_{1}+(n_{2}^{4}-4n_{2}^{3}+3n_{2}^{2}+4n_{2}-4)\Big{)};\\ &z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-3,3}^{\circ}\big{(}(n_{1},n_{2})\big{)}=\frac{n_{1}n_{2}}{12}\Big{(}n_{1}^{4}+2(n_{2}-3)n_{1}^{3}+(4n_{2}^{2}-12n_{2}+13)n_{1}^{2}\\ &\qquad\qquad+2(n_{2}^{3}-6n_{2}^{2}+10n_{2}-6)n_{1}+(n_{2}^{4}-6n_{2}^{3}+13n_{2}^{2}-12n_{2}+4)\Big{)};\\ &z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-4,2}^{\circ}\big{(}(n_{1},n_{2})\big{)}=\frac{n_{1}n_{2}}{72}\Big{(}n_{1}^{6}+3(n_{2}-3)n_{1}^{5}+2(3n_{2}^{2}-12n_{2}+14)n_{1}^{4}\\ &\qquad\qquad+\frac{3}{2}(5n_{2}^{3}-26n_{2}^{2}+41n_{2}-20)n_{1}^{3}+(6n_{2}^{4}-39n_{2}^{3}+78n_{2}^{2}-45n_{2}-11)n_{1}^{2}\\ &\qquad\qquad+\frac{3}{2}(2n_{2}^{5}-16n_{2}^{4}+41n_{2}^{3}-30n_{2}^{2}-19n_{2}+26)n_{1}\\ &\qquad\qquad+(n_{2}^{6}-9n_{2}^{5}+28n_{2}^{4}-30n_{2}^{3}-11n_{2}^{2}+39n_{2}-18)\Big{)}.\end{split}$
Example 5.6.
Similarly, we give some conjectural formulas of the two-point connected correlators for $r=3$. Let $n_{1}\geq n_{2}>0$ and denote $n=n_{1}+n_{2}$. First by the Riemann-Hurwitz formula, we have $N_{n-k_{1},n-k_{2},k_{3}}^{\circ}\big{(}(n_{1},n_{2})\big{)}=0$ unless $k_{1}+k_{2}-k_{3}$ is a nonnegative even number. Using the numerical data obtained from the results in Example 3.6, we are able to check the following formulas for small $n_{1},n_{2}$:
$\begin{split}&z_{(n_{1},n_{2})}\cdot N_{n-1,n-1,2}^{\circ}\big{(}(n_{1},n_{2})\big{)}=n_{1}n_{2}\Big{(}n_{1}^{2}+(n_{2}-1)n_{1}+(n_{2}^{2}-n_{2})\Big{)},\\ &z_{(n_{1},n_{2})}\cdot N_{n-2,n-1,1}^{\circ}\big{(}(n_{1},n_{2})\big{)}=\frac{n_{1}n_{2}}{24}\Big{(}3n_{1}^{4}+2(3n_{2}-2)n_{1}^{3}+3(2n_{2}^{2}-2n_{2}-1)n_{1}^{2}\\ &\qquad\qquad+2(3n_{2}^{3}-3n_{2}^{2}-3n_{2}+2)n_{1}+(3n_{2}^{4}-4n_{2}^{3}-3n_{2}^{2}+4n_{2})\Big{)}.\end{split}$
Remark 5.1.
In Examples 5.2-5.3 we have seen that the $1$-point correlators for small $r$ and $k_{i}$ can be represented as a linear combination of the binomial coefficients $\binom{n}{d}$ where the coefficients are nonnegative integers. One can check that this property also holds for the conjectural formulas in Examples 5.5-5.6. For instance,
$\begin{split}z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-1,1}^{\circ}\big{(}(n_{1},n_{2})\big{)}=&\binom{n_{1}}{1}\binom{n_{2}}{1},\\ z_{(n_{1},n_{2})}\cdot N_{n_{1}+n_{2}-2,2}^{\circ}\big{(}(n_{1},n_{2})\big{)}=&3\binom{n_{1}}{3}\binom{n_{2}}{1}+3\binom{n_{1}}{1}\binom{n_{2}}{3}+2\binom{n_{1}}{2}\binom{n_{2}}{2}\\ &+2\binom{n_{1}}{2}\binom{n_{2}}{1}+2\binom{n_{1}}{1}\binom{n_{2}}{2}.\end{split}$
It is natural to make another conjecture that such a property holds for all connected correlators $z_{\mu}\cdot N_{|\mu|-k_{1},|\mu|-k_{2},\cdots,|\mu|-k_{r-1},k_{r}}^{\circ}(\mu)$. We call it the conjectural non-negative integral coefficient binomial expansion. We will study this interesting phenomenon in future works.
5.2. Proof of the polynomiality for $1$-point correlators
In this subsection we prove Conjecture 5.1 in the one-point case:
Proposition 5.1.
For every fixed $r$ and $k_{1},k_{2},\cdots,k_{r}$, there exists a polynomial $A_{k_{1},\cdots,k_{r}}(x)\in{\mathbb{Q}}[x]$ such that
(65) $n\cdot N_{n-k_{1},\cdots,n-k_{r-1},k_{r}}^{\circ}\big{(}(n)\big{)}=A_{k_{1},\cdots,k_{r}}(n)$
for every integer $n$ satisfying $n\geq\max\{k_{1},\cdots,k_{r}\}$.
Notice that the correlator $n\cdot N_{n-k_{1},\cdots,n-k_{r-1},k_{r}}^{\circ}\big{(}(n)\big{)}$ is a rational number for every $n$, and thus if there exists a polynomial satisfying (65), it must have rational coefficients. In what follows, we prove the existence of the polynomial satisfying (65). By (61), this is equivalent to saying that
(66) $\begin{split}&\frac{1}{n}\cdot\sum_{a=0}^{n-1}\Big{[}\Big{(}\prod_{j=1}^{r-1}\sum_{-a\leq i_{1}<\cdots<i_{k_{j}}\leq n-1-a}i_{1}\cdots i_{k_{j}}\Big{)}\\ &\cdot\Big{(}\sum_{j=0}^{k_{r}-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k_{r}-1-j}^{\prime}\leq n-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k_{r}-1-j}^{\prime}}\Big{)}\Big{]}\end{split}$
is a polynomial in $n$. We prove this in several steps. First we show that:
Lemma 5.1.
For the fixed positive integer $k$,
$f_{k}(n,a)=\sum_{-a\leq i_{1}<\cdots<i_{k}\leq n-1-a}i_{1}\cdots i_{k}$
is a polynomial in $n$ and $a$.
Proof.
This can be proved by induction on $k$. First for $k=1$,
$f_{1}(n,a)=\sum_{-a\leq i\leq n-1-a}i=\frac{1}{2}n(n-2a-1)$
is a polynomial in $n$ and $a$. Now assume $f_{k}(n,a)$ is a polynomial, then
$\begin{split}f_{k+1}(n,a)=&\sum_{i_{k}=-a+k}^{n-1-a}i_{k}\cdot\Big{(}\sum_{-a\leq i_{1}<\cdots<i_{k}\leq i_{k}-1}i_{1}\cdots i_{k}\Big{)}\\ =&(-a+k)f_{k}(-a+k-1,a)+\cdots+(n-1-a)f_{k}(n-2-a,a)\end{split}$
is also a polynomial in $n$ and $a$. ∎
Then (66) becomes:
(67) $\frac{1}{n}\sum_{a=0}^{n-1}\Big{(}\prod_{j=1}^{r-1}f_{k_{j}}(n,a)\Big{)}\Big{(}\sum_{j=0}^{k_{r}-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k_{r}-1-j}^{\prime}\leq n-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k_{r}-1-j}^{\prime}}\Big{)},$
Now we can prove the special case $k_{r}=1$ of Proposition 5.1:
Lemma 5.2.
For every fixed $r$ and $k_{1},k_{2},\cdots,k_{r-1}$, the one-point correlator
$n\cdot N_{n-k_{1},\cdots,n-k_{r-1},1}^{\circ}\big{(}(n)\big{)}=\frac{1}{n}\sum_{a=0}^{n-1}\Big{(}\prod_{j=1}^{r-1}f_{k_{j}}(n,a)\Big{)}$
is a polynomial in $n$.
Proof.
by Lemma 5.1 we see that $\prod_{j=1}^{r-1}f_{k_{j}}(n,a)$ is a polynomial in $n$ and $a$. Denote:
$\prod_{j=1}^{r-1}f_{k_{j}}(n,a)=g_{0}(n)+g_{1}(n)a+\cdots+g_{M}(n)a^{M},$
where $g_{i}(n)$ are polynomials in $n$, and then:
$\frac{1}{n}\sum_{a=0}^{n-1}\Big{(}\prod_{j=1}^{r-1}f_{k_{j}}(n,a)\Big{)}=\sum_{j=0}^{M}g_{j}(n)\cdot\frac{1}{n}\sum_{a=0}^{n-1}a^{j}.$
It is well-known that $\frac{1}{n}\sum_{a=0}^{n-1}a^{j}$ is a polynomial in $n$ for every $j\geq 0$. ∎
To prove the general case of Proposition 5.1, we need the following:
Lemma 5.3.
Assume $k>1$, then:
(68) $\sum_{a=0}^{n-1}\Big{(}\sum_{j=0}^{k-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-1-j}^{\prime}\leq n-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-1-j}^{\prime}}\Big{)}=0.$
Proof.
By (61), this is equivalent to saying that $N_{k}^{\circ}\big{(}(n)\big{)}=0$ for every $k>1$. This has been proved by the Riemann-Hurwitz formula in Example 5.1. ∎
The following corollary will also be useful in the proof of Proposition 5.1:
Corollary 5.1.
Let $n$, $k$ be two integers satisfying $n\geq k>1$. Then:
(69) $\begin{split}&\sum_{-n\leq i_{1}<\cdots<i_{k-1}\leq-1}\frac{1}{i_{1}\cdots i_{k-1}}\\ &\qquad+\sum_{a=0}^{n-1}\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq n-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}(n-a)}=0.\end{split}$
Here for $k=2$ we use the convention:
$\sum_{j=0}^{0}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq n-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}(n-a)}=\frac{1}{n-a}.$
Proof.
Denote by $R(n,k)$ the left-hand side of (68), then the left-hand side of (69) is actually $R(n+1,k)-R(n,k)$ and thus equals to zero by Lemma 5.3. ∎
Now we are able to prove Proposition 5.1. The special case $k_{r}=1$ has been proved in Lemma 5.2, and thus we only need to prove for $k_{r}>1$. Recall that by Lemma 5.1, the term
$\Big{(}\prod_{j=1}^{r-1}\sum_{-a\leq i_{1}<\cdots<i_{k_{j}}\leq n-1-a}i_{1}\cdots i_{k_{j}}\Big{)}$
in (66) is a polynomial in $n$ and $a$, whose expression depends only on the numbers $k_{1},\cdots,k_{r-1}$. Therefore it is sufficient to show the following:
Lemma 5.4.
Let $f(x,a)$ be a polynomial in $x$ and $a$. Then for every integer $k>1$, there exists a polynomial $P_{f,k}(x)$ in $x$ such that
(70) $P_{f,k}(x)=\sum_{a=0}^{n-1}f(n,a)\cdot\Big{(}\sum_{j=0}^{k-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-1-j}^{\prime}\leq n-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-1-j}^{\prime}}\Big{)}$
for every positive integer $n$, and $P_{f,k}(x)$ can be divided by $x$.
Proof.
Assume that
$f(x,a)=g_{0}(a)+g_{1}(a)x+\cdots+g_{m}(a)x^{m}.$
Denote by ${\widetilde{P}}_{f,k}(n)$ the right-hand side of (70), then:
$\begin{split}{\widetilde{P}}_{f,k}(n)=&\sum_{a=0}^{n-1}g_{0}(a)\Big{(}\sum_{j=0}^{k-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-1-j}^{\prime}\leq n-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-1-j}^{\prime}}\Big{)}\\ &+\sum_{s=1}^{m}n^{s}\Big{[}\sum_{a=0}^{n-1}g_{s}(a)\Big{(}\sum_{j=0}^{k-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-1-j}^{\prime}\leq n-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-1-j}^{\prime}}\Big{)}\Big{]},\end{split}$
and therefore it suffices to show that for every polynomial $g(a)$ and $k>1$, there exists a polynomial $Q_{g,k}(x)$ such that
(71) $Q_{g,k}(n)=\sum_{a=0}^{n-1}g(a)\Big{(}\sum_{j=0}^{k-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-1-j}^{\prime}\leq n-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-1-j}^{\prime}}\Big{)}$
for every positive integer $n$, and $Q_{g,k}(x)$ is divided by $x$. In what follows, we prove this by induction on the degree of $g(a)$.
Denote by ${\widetilde{Q}}_{g,k}(n)$ the right-hand side of (71). If $\deg(g)=0$, i.e., $g(a)$ is a constant, then by Lemma 5.3 we know that ${\widetilde{Q}}_{g,k}(n)=0$ for every $n\geq 1$. In this case one simply takes ${\widetilde{Q}}_{g,k}(n)=0$.
Now assume the conclusion holds for all $g(a)$ satisfying $\deg(g)\leq d$, and let ${\tilde{g}}(a)$ be a polynomial of degree $d+1$. Let $b\geq 1$ be an integer, then for every integer $b\geq 1$, one has:
$\begin{split}&{\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)={\tilde{g}}(b)\cdot\sum_{-b\leq i_{1}<\cdots<i_{k-1}\leq-1}\frac{1}{i_{1}\cdots i_{k-1}}\\ &\qquad\quad+\sum_{a=0}^{b-1}{\tilde{g}}(a)\cdot\Big{(}\sum_{j=0}^{k-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-1-j}^{\prime}\leq b-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-1-j}^{\prime}}\Big{)}\\ &\qquad\quad-\sum_{a=0}^{b-1}{\tilde{g}}(a)\cdot\Big{(}\sum_{j=0}^{k-1}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-1-j}^{\prime}\leq b-1-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-1-j}^{\prime}}\Big{)},\end{split}$
and thus:
$\begin{split}&{\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)={\tilde{g}}(b)\cdot\sum_{-b\leq i_{1}<\cdots<i_{k-1}\leq-1}\frac{1}{i_{1}\cdots i_{k-1}}\\ &\qquad+\sum_{a=0}^{b-1}{\tilde{g}}(a)\cdot\Big{(}\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq b-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}\cdot(b-a)}\Big{)}.\end{split}$
Notice that ${\tilde{g}}(a)-{\tilde{g}}(b)$ is a polynomial in $b$ and $a$ which is divided by $(b-a)$. Define a polynomial $h(b,a)$ by:
$h(b,a)=\frac{{\tilde{g}}(a)-{\tilde{g}}(b)}{b-a},$
then ${\tilde{g}}(a)=(b-a)\cdot h(b,a)+{\tilde{g}}(b)$, and thus:
(72) $\begin{split}&{\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)={\tilde{g}}(b)\cdot\Big{(}\sum_{-b\leq i_{1}<\cdots<i_{k-1}\leq-1}\frac{1}{i_{1}\cdots i_{k-1}}\\ &\qquad+\sum_{a=0}^{b-1}\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq b-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}\cdot(b-a)}\Big{)}\\ &\qquad+\sum_{a=0}^{b-1}h(b,a)\Big{(}\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq b-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}}\Big{)}\end{split}$
(Notice here if $k=2$, the term
$\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq n+c-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}}$
in the third line of (72) should be understood as $1$). Now use Corollary 5.1, and then we obtain:
$\begin{split}&{\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)\\ =&\sum_{a=0}^{b-1}h(b,a)\Big{(}\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq b-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}}\Big{)}.\end{split}$
Assume $h(b,a)=h_{0}(a)+h_{1}(a)\cdot b+\cdots h_{M}(a)\cdot b^{M}$, then:
$\begin{split}&{\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)\\ =&\sum_{s=0}^{M}b^{s}\sum_{a=0}^{b-1}h_{s}(a)\Big{(}\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq b-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}}\Big{)}.\end{split}$
Here by the definition of $h(n,a)$ we know that $h_{s}(a)$ is a polynomial in $a$ of degree less than $d+1$. Thus if $k>2$, by the induction hypothesis there exists a polynomial $q_{s}(x)$ such that $x$ divides $q_{s}(x)$ and
$q_{s}(b)=\sum_{a=0}^{b-1}h_{s}(a)\Big{(}\sum_{j=0}^{k-2}\sum_{\begin{subarray}{c}-a\leq i_{1}<\cdots<i_{j}\leq-1\\ 1\leq i_{1}^{\prime}<\cdots<i_{k-2-j}^{\prime}\leq b-a\end{subarray}}\frac{1}{i_{1}\cdots i_{j}\cdot i_{1}^{\prime}\cdots i_{k-2-j}^{\prime}}\Big{)}$
for every positive integer $b$. Now denote $\tilde{q}(x)=\sum_{s=0}^{M}x^{s}\cdot q_{s}(x)$, then $x$ divides $\tilde{q}(x)$ and ${\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)=\tilde{q}(b)$ for every positive integer $b$.
For the special case $k=2$, we have
${\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)=\sum_{a=0}^{b-1}h(b,a)=\sum_{s=0}^{M}b^{s}\sum_{a=0}^{b-1}h_{s}(a).$
It is well-known that for every integer $j\geq 0$, the summation $\sum_{i=0}^{b-1}i^{j}$ is a polynomial in $b$ which can be divided by $b$, and thus the summation $\sum_{a=0}^{b-1}h_{s}(a)$ is a polynomial in $b$ which is divided by $b$. We denote this polynomial by $\tilde{q}(x)$.
In either case, we obtain a polynomial $\tilde{q}(x)$ such that $x$ divides $\tilde{q}(x)$ and
$\tilde{q}(b)={\widetilde{Q}}_{{\tilde{g}},k}(b+1)-{\widetilde{Q}}_{{\tilde{g}},k}(b)$
for every positive integer $b$. Now we have:
$\begin{split}{\widetilde{Q}}_{{\tilde{g}},k}(n)=&\sum_{b=1}^{n-1}\Big{(}{\widetilde{Q}}_{{\tilde{g}},k}(j+1)-{\widetilde{Q}}_{{\tilde{g}},k}(j))\Big{)}+{\widetilde{Q}}_{{\tilde{g}},k}(1)\\ =&\sum_{b=1}^{n-1}\tilde{q}(b)+{\widetilde{Q}}_{{\tilde{g}},k}(1).\end{split}$
Here ${\widetilde{Q}}_{{\tilde{g}},k}(1)=0$ for every $k>1$, and $\tilde{q}(x)$ is of the form $\tilde{q}(x)=\sum_{j=1}^{J}\tilde{q}_{j}\cdot x^{j}$ since $x$ divides $\tilde{q}(x)$. Notice that for every $j\geq 1$, the summation $\sum_{i=1}^{n-1}i^{j}$ is a polynomial in $n$ which can be divided by $n$, and therefore there exists a polynomial $Q_{{\tilde{g}},k}(x)$ such that $x$ divides $Q_{{\tilde{g}},k}(x)$ and
$Q_{{\tilde{g}},k}(n)=\sum_{b=1}^{n-1}\tilde{q}(b)={\widetilde{Q}}_{{\tilde{g}},k}(n)$
for every positive integer $n$. This completes the proof. ∎
Acknowledgements. The authors would like to thank Professor Xiaobo Liu and Professor Jian Zhou for helpful discussions and encouragement. The second-named author is supported by the NSFC grant (No. 12288201), the China Postdoctoral Science Foundation (No. 2023M743717), and the China National Postdoctoral Program for Innovative Talents (No. BX20240407).
Appendix A A Proof of the Evaluation Formula for $s_{\mu}(t_{k}=\frac{v}{k})$ via $W$-Operators
In this appendix, we give a new proof of Lemma 3.1. Our strategy is to use the action of $W$-operators $P_{m}^{(k)}$ on Schur functions.
First we take $m=\pm 1$ and $k=1,2$ in Theorem 3.1, and obtain:
(73) $t_{1}\cdot s_{\mu}(\bm{t})=\sum_{\mu+\square\in{\mathcal{P}}}s_{\mu+\square}(\bm{t}),\qquad L_{-1}s_{\mu}(\bm{t})=\sum_{\mu+\square\in{\mathcal{P}}}c(\square)\cdot s_{\mu+\square}(\bm{t}),$
and
(74) $\frac{\partial}{\partial t_{1}}s_{\mu}(\bm{t})=\sum_{\mu-\square\in{\mathcal{P}}}s_{\mu-\square}(\bm{t}),\qquad L_{1}s_{\mu}(\bm{t})=\sum_{\mu-\square\in{\mathcal{P}}}c(\square)\cdot s_{\mu-\square}(\bm{t}),$
where
$\begin{split}L_{-1}=\sum_{n\geq 1}(n+1)t_{n+1}\frac{\partial}{\partial t_{n}},\qquad\qquad L_{1}=\sum_{n\geq 1}nt_{n}\frac{\partial}{\partial t_{n+1}}\end{split}$
are the Virasoro operators. Here we denote by $\mu-\square$ a partition whose Young diagram is obtained from that of $\mu$ by deleting a box $\square$ on the right of some row, and set $s_{\mu-\square}=0$ if $\mu-\square$ is no longer a partition. The above formulas for $t_{n}s_{\mu}$ and $\frac{\partial}{\partial t_{n}}s_{\lambda}$ are well-known in literatures, and formulas for $k=2$ can also be found in [1, 28]. Now let $v$ be a formal variable, and we prove the following:
Lemma A.1.
We have:
(75) $e^{s(L_{-1}+v\cdot t_{1})}(1)=\sum_{\mu\in{\mathcal{P}}}s^{|\mu|}\cdot s_{\mu}(t_{k}=\frac{v}{k})\cdot s_{\mu}(\bm{t}).$
Proof.
Denote by $G(\bm{t})$ the right-hand side of (75). By (73) we have:
$\begin{split}(L_{-1}+vt_{1})G(\bm{t})=&\sum_{\mu\in{\mathcal{P}}}s^{|\mu|}\cdot s_{\mu}(t_{k}=\frac{v}{k})\cdot\sum_{\mu+\square\in{\mathcal{P}}}\big{(}c(\square)+v\big{)}s_{\mu+\square}(\bm{t})\\ =&\sum_{\lambda\in{\mathcal{P}}}s^{|\lambda|-1}\cdot\sum_{\lambda-\square\in{\mathcal{P}}}\big{(}c(\square)+v\big{)}s_{\lambda-\square}(t_{k}=\frac{v}{k})\cdot s_{\lambda}(\bm{t}),\end{split}$
where we denote $\lambda=\mu+\square$. Then by (74) we see that:
$\begin{split}(L_{-1}+vt_{1})G(\bm{t})=&\sum_{\lambda\in{\mathcal{P}}}s^{|\lambda|-1}\cdot(L_{1}+v\frac{\partial}{\partial t_{1}})s_{\lambda}(\bm{t})\big{|}_{t_{k}=\frac{v}{k}}\cdot s_{\lambda}(\bm{t})\\ =&\sum_{\lambda\in{\mathcal{P}}}s^{|\lambda|-1}\cdot\Big{(}v\frac{\partial}{\partial t_{1}}+\sum_{n\geq 1}p_{n}\cdot(n+1)\frac{\partial}{\partial p_{n+1}}\Big{)}s_{\lambda}(\bm{t})\big{|}_{p_{k}=v}\cdot s_{\lambda}(\bm{t})\end{split}$
where $p_{k}=kt_{k}$ for every $k$. Recall that if we set $\deg(t_{k})=\deg(p_{k})=k$, then $s_{\lambda}(\bm{t})$ is a weighted homogeneous polynomial of degree $|\mu|$. Thus by Euler’s formula,
$\Big{(}v\frac{\partial}{\partial t_{1}}+\sum_{n\geq 1}p_{n}\cdot(n+1)\frac{\partial}{\partial p_{n+1}}\Big{)}s_{\lambda}(\bm{t})\big{|}_{p_{k}=v}=|\lambda|\cdot s_{\lambda}(\bm{t})\big{|}_{p_{k}=v}=|\lambda|\cdot s_{\lambda}(t_{k}=\frac{v}{k}),$
and then:
$\begin{split}(L_{-1}+vt_{1})G(\bm{t})=\sum_{\lambda\in{\mathcal{P}}}s^{|\lambda|-1}\cdot|\lambda|\cdot s_{\lambda}(t_{k}=\frac{v}{k})\cdot s_{\lambda}(\bm{t})=\frac{\partial}{\partial s}G(t).\end{split}$
Notice that $G(\bm{t})|_{s=0}=1$, and then we conclude that $G(\bm{t})=e^{s(L_{-1}+v\cdot t_{1})}(1)$. ∎
Now we can present a proof of Lemma 3.1. We expand the exponential in the left-hand side of (75) directly and then apply (73)
$\begin{split}e^{s(L_{-1}+v\cdot t_{1})}(1)=&\sum_{n=0}^{\infty}\frac{s^{n}}{n!}\cdot(L_{-1}+v\cdot t_{1})^{n}s_{(\emptyset)}(\bm{t})\\ =&\sum_{n=0}^{\infty}\frac{s^{n}}{n!}\cdot\sum_{|\mu|=n}K_{\mu}\cdot\prod_{\square\in\mu}\big{(}c(\square)+v\big{)}s_{\mu}(\bm{t}),\end{split}$
where $K_{\mu}$ is the number of standard Young tableaux of shape $\mu$. By (36) we have:
(76) $e^{s(L_{-1}+v\cdot t_{1})}(1)=\sum_{\mu\in{\mathcal{P}}}s^{|\mu|}\prod_{\square\in\mu}\frac{\big{(}c(\square)+v\big{)}}{h(\square)}s_{\mu}(\bm{t}).$
Now compare this equality with (75), and we obtain Lemma 3.1.